### v3

# Generalized weight properties of resultants and discriminants, and applications to projective enumerative geometry

Generalized weight properties of resultants and discriminants, and applications to projective enumerative geometry
by Laurent Busé and Thomas Dedieu footnotetext: This text is from a series originated in the seminar Degenerations and enumeration of curves on surfaces, held at the University of Roma “Tor Vergata” in the years 2015–2017. See https://www.math.univ-toulouse.fr/˜tdedieu/#enumTV for the full collection.
In his wonderful book [18], first published in 1862, Salmon casually gives the leading term of some Taylor expansion of the discriminant of a plane algebraic curve. Consider the curve $\mathrm{V}(F)\subseteq\mathbf{P}^{2}$; in a suitable system of homogeneous coordinates (X:Y:Z)(X\!:\!Y\!:\!Z), its equation $F$ has the form
$F(X,Y,Z)=TZ^{d-1}Y+\frac{1}{2}Z^{d-2}(AX^{2}+2BXY+CY^{2})+\sum_{k\geqslant 3}Z^{d-k}F_{k}(X,Y)=0,$
where each polynomial $F_{k}(X,Y)$ is homogeneous of degree $k$ in $X,Y$, and $T,A,B,C$ are coefficients of $F$. Salmon states, without any explanation or reference, that the discriminant of $F$ is of the form
[…]
where $\Phi$ is “the discriminant when $T$ vanishes” (and $\Psi$ is some polynomial in the coefficients of $F$) [18, § 605]. This note arose as an attempt to understand and prove this formula.
We introduce the reduced discriminant of a hypersurface (or homogeneous polynomial), of which $\Phi$ is an instance, it turns out. It is deeply linked with the reduced resultant, introduced by Zariski in [21], in 1936 long after Salmon. Of course, when $T$ is zero, $\mathrm{V}(F)$ is singular at the point $(0:0:1)$, no matter the other coefficients of $F$, and correspondingly $\mathrm{Disc}(F)$ vanishes identically. The polynomial $\Phi$, in fact, vanishes at those values of the coefficients of $F$ for which the curve $\mathrm{V}(F)$ has some additional singularity besides the point $(0:0:1)$. The reduced discriminant generalizes this to the case when $F$ has a singular point of arbitrary multiplicity $s$. Somehow built in the theory is a general form of Formula (a). We show how to get (a) specifically (in fact with a normalization factor $\frac{1}{2}$, which is probably irrelevant for Salmon), and generalize the formula for hypersurfaces of arbitrary dimension, see Section 2.3. This is certainly what Salmon has in mind, although arguably only for $s=2$, and possibly only empirically. It has since then been apparently completely forgotten; we have only found a faint trace of these ideas in Salmon’s works, see Paragraph (2.12). We point out that the reduced discriminant is a particular instance of the theory of toric discriminants studied in [12], which is more general in terms of the family of hypersurfaces that can be considered. The methodology that we develop here in the specific setting of the reduced discriminant is different from the theory of toric discriminants, and we found it particularly well adapted, if not necessary, to tackle Salmon’s formula.
Salmon then uses Formula (a) to derive various enumerative quantities for surfaces $S\subseteq\mathbf{P}^{3}$. In particular, he computes the number of bitangent planes passing through a fixed general point $p\in\mathbf{P}^{3}$. We explain his method, which involves a remarkable trick in elimination theory, and generalize it for hypersurfaces of arbitrary dimension. Salmon’s strategy is to consider a pencil of planes with center a line tangent to the surface $S$ at some point $p^{\prime}$; this pencil contains a finite number of planes tangent to $S$, among which the tangent plane at $p^{\prime}$ counts with multiplicity $2$ in general, and with greater multiplicity if it has some special feature, e.g., if it is a bitangent plane. In an appropriate setting, this multiplicity is the order in $T$ of the polynomial in (a), and the game is to understand the conditions on the point $p^{\prime}$ that make this multiplicity jump. In the upshot we get, indirectly from the reduced discriminant $\Phi$, a curve in $S$ parametrizing those points $p^{\prime}$ such that the tangent plane at $p^{\prime}$ is a bitangent plane; following Salmon, we call it the node-couple curve. There are other well-known ways to compute the number of bitangent planes to a surface in $\mathbf{P}^{3}$ (see, e.g., [II]). The present method has the advantage of being core level: it is only a study of homogeneous polynomials and their discriminants.
The present text includes two other projective enumerative computations taken from Salmon’s book; they both fit in the framework of what we would call reduced elimination theory. The first one gives the number of bitangent lines to a surface in $\mathbf{P}^{3}$ through a general point. We generalize it in two ways: to the enumeration of lines in $\mathbf{P}^{n}$ having contact of the form $2p+(n-1)q+\cdots$ with a hypersurface on the one hand, and to the enumeration of bitangent codimension 2 linear spaces on the other hand. While Salmon’s computation and the first generalization do not involve openly the reduced discriminant (thanks to the existence of a cheap substitute in the two indeterminates case), it is essential for the latter generalization. The second computation is that of the degree of the flecnodal curve of a surface $S\subseteq\mathbf{P}^{3}$, which is the locus of points $p$ at which the tangent plane cuts out a curve with one of its two branches at $p$ that has an inflection point (equivalently, there is a line having contact order with $S$ at $p$ at least $4$). We present here Salmon’s ideas, noting that this computation is carried out in modern standards and in arbitrary dimension in [3].
We also give a synthetic account of the theory of resultants and discriminants, in its modern form worked out by Jouanolou, see, e.g., [15]. On the one hand this is necessary for us to develop the theory of the reduced discriminant (the reduced resultant had already been treated by Ould Mohamdi in [16], under the direction of Jouanolou; note, however, that passing from the resultant to the discriminant is not straightforward, see, e.g., , the comments after Definition (2.3)). On the other hand, we believe this text is a good occasion to make Jouanolou’s formalism accessible to the XXIst century classical algebraic geometer.
We put particular focus on the homogeneity properties of the resultants and discriminants, with respect to various gradings. They are indeed essential for the enumerative applications, and we give numerous illustrations, including a proof of Bezout’s theorem and the computation of the degree of the dual of a smooth hypersurface. Beyond that, they are central in the theory of reduced resultants and discriminants, see Theorem (1.3) and Corollary (2.13), and indeed the key to Formula (a) and its generalizations.
ThD thanks Ragni Piene and Israel Vainsencher for their interest, and for suggesting he talk with LB; he also thanks François Lê for sharing some insights on Salmon’s work. LB is grateful to Alexandru Dimca for useful discussions on Milnor number and for indicating the reference [11]. We both thank the referees for their helpful observations.
(0.1)Notation
Let $\mathbf{k}$ be a commutative ring, $n\geqslant 1$ be an integer, and $\mathbf{X}:=(X_{0},X_{1},\ldots,X_{n})$ be a sequence of indeterminates. Given a multi-index ${\boldsymbol{\alpha}}:=(\alpha_{0},\alpha_{1},\ldots,\alpha_{n})\in\mathbf{N}^{n}$, we denote by $\mathbf{X}^{\boldsymbol{\alpha}}$ the monomial $X_{0}^{\alpha_{0}}X_{1}^{\alpha_{1}}\ldots X_{n}^{\alpha_{n}}$, and set $|{\boldsymbol{\alpha}}|=\sum_{i=0}^{n}\alpha_{i}$.
(0.2) Definition.
Given a grading $\mathbf{k}[\mathbf{X}]=\bigoplus_{d\geqslant 0}\mathbf{k}[\mathbf{X}]_{d}$, the order of a polynomial $F\in\mathbf{k}[\mathbf{X}]$ is the largest integer $e$ such that $F\in\bigoplus_{d\geqslant e}\mathbf{k}[\mathbf{X}]_{d}$.
We will usually consider gradings obtained by assigning a weight $w_{i}$ to each indeterminate $X_{i}$: the degree $d$ homogeneous piece $\mathbf{k}[\mathbf{X}]_{d}$ is then generated by all monomials $\mathbf{X}^{\boldsymbol{\alpha}}$ such that $\sum_{i=0}^{n}\alpha_{i}w_{i}=d$.
Given a subset $I\subseteq\{1,\ldots,n\}$, the order of a polynomial $F$ in the indeterminates $X_{i}$, $i\in I$, is the order of $F$ with respect to the grading obtained by assigning to $X_{i}$ the weight $1$ if $i\in I$, and the weight 0 if $i\not\in I$.
Beware that $\mathbf{k}$ itself will often be a polynomial ring, and we will consider gradings and orders both on $\mathbf{k}[\mathbf{X}]$ and on $\mathbf{k}$ itself.
Reduced resultant
Let $F_{1},\ldots,F_{n+1}$ be $n+1$ homogeneous polynomials in the indeterminates $X_{0},X_{1},\ldots,X_{n}$. They define a collection of $n+1$ hypersurfaces in the projective space $\mathbf{P}^{n}$, and the intersection of these $n+1$ hypersurfaces is empty if $F_{1},\ldots,F_{n+1}$ are general. In fact, the emptiness of this intersection is characterized by the non-vanishing of the resultant $\mathrm{Res}(F_{1},\ldots,F_{n+1})$ of these polynomials (see §1.1). Thus, the resultant characterizes those collections of homogeneous polynomials that have a common non-trivial root. The purpose of the reduced resultant is similar: one still considers $n+1$ homogeneous polynomials $F_{1},\ldots,F_{n+1}$, but this time one assumes that they have a common non-trivial root; then the reduced resultant will characterize those collections of polynomials having an additional common non-trivial root. It has been introduced by Zariski [21]; a more complete and modern treatment is given in [16].
In what follows, we use the resultant of multivariate homogeneous polynomials as developed in [15] (see also [7, 8, 12]). An introduction and a brief recap are given in Section 1.1 below.
Inertia forms and the ordinary multivariate resultant
We follow the beautiful presentation in [20, Chapter XI] (beware that this Chapter on Elimination Theory has disappeared in later editions).
(1.1)Saturation of a homogeneous ideal
We recall the following for the convenience of the reader; see, e.g., [13, Lecture 5] or [14, Exercise II.5.10] for more details. The saturation of a homogeneous ideal $I\subseteq\mathbf{k}[\mathbf{X}]$ is the homogeneous ideal
(1.1.1) $\bar{I}$ $=\bigl\{F\in\mathbf{k}[\mathbf{X}]:\forall i=0,\ldots,n,\ \exists N_{i}\ \text{such that}\ X_{i}^{N_{i}}F\in I\bigr\}$
$=I:(X_{0},\ldots,X_{n})^{\infty}.$
For sufficiently large $m$, the graded pieces $I_{m}$ and $\bar{I}_{m}$ are equal. Moreover, for two homogeneous ideals $I$ and $J$, the following three propositions are equivalent:
  1. (i)
    $\bar{I}=\bar{J}$;
  2. (ii)
    $I_{m}=J_{m}$ for sufficiently large $m$;
  3. (iii)
    $I\cdot\mathbf{k}[\mathbf{X},\frac{1}{X_{i}}]=J\cdot\mathbf{k}[\mathbf{X},\frac{1}{X_{i}}]$ for all $i=0,\ldots,n$.
In other words, a subscheme $V\subseteq\mathbf{P}^{n}_{\mathbf{k}}$ is defined (scheme-theoretically) by a homogeneous ideal $I\subseteq\mathbf{k}[\mathbf{X}]$ if and only if the saturation $\bar{I}$ equals the homogeneous ideal $I_{V}$ of $V$. For example, for all $F\in\mathbf{k}[\mathbf{X}]$, the two ideals $(F)$ and $(X_{0}F,\ldots,X_{n}F)$ define the same subcheme of $\mathbf{P}^{n}_{\mathbf{k}}$ (and we recommend that the reader check it for themselves).
In particular, if $\mathbf{k}$ is a field, a subscheme $V\subseteq\mathbf{P}^{n}_{\mathbf{k}}$ defined by a homogeneous ideal $I$ is empty if and only if the degree 0 piece $\bar{I}_{0}$ is non-zero: this is the homogeneous Nullstellensatz! (see also [6]). On the other hand, the non-emptiness of $V$ is equivalent to the existence of a point in $V(\mathbf{k}^{\prime})$ for some finite field extension $\mathbf{k}^{\prime}$ of $\mathbf{k}$.
When $\mathbf{k}$ is an arbitrary commutative integral domain,111this arguably unconventional notation is used ubiquitously by Jouanolou. the vanishing of $\bar{I}_{0}$ is equivalent to the scheme-theoretic image of the map $V\to\mathrm{Spec}(\mathbf{k})$ being equal to the whole $\mathrm{Spec}(\mathbf{k})$. Indeed, the subscheme of $\mathrm{Spec}(\mathbf{k})$ defined by $\bar{I}_{0}$ coincides as a set with the image of $V\to\mathrm{Spec}(\mathbf{k})$ — this is the proof that projective morphisms are closed —, and moreover $\bar{I}_{0}$ defines the scheme-theoretic image of $V\to\mathrm{Spec}(\mathbf{k})$ (see, e.g., [15, §1] for more details).
Let $d_{1},\ldots,d_{r}$ be positive integers. For all $j=1,\ldots,r$, we consider the generic homogeneous degree $d_{j}$ polynomial in the indeterminates $\mathbf{X}=(X_{0},\ldots,X_{n})$
$F_{j}:=\sum_{|{\boldsymbol{\alpha}}|=d_{j}}U_{j,{\boldsymbol{\alpha}}}\mathbf{X}^{\boldsymbol{\alpha}}.$
“Generic” refers to the fact that the coefficients of the $F_{j}$’s are indeterminates; indeed, $F_{j}$ is an incarnation of the generic point, in the sense of schemes, of the affine space parametrizing homogeneous degree $d_{j}$ polynomials.
We let $A_{\mathbf{Z}}$ be the algebra generated by the indeterminates coefficients of the $F_{j}$’s, i.e., we set $A_{\mathbf{Z}}:=\mathbf{Z}\bigl[U_{j,{\boldsymbol{\alpha}}}\bigr]_{j=1,\ldots,r,\,|{\boldsymbol{\alpha}}|=d_{j}}$. Thus, $F_{j}\in A_{\mathbf{Z}}{}[\mathbf{X}]$ for all $j=1,\ldots,r$.
Let $\mathbf{k}$ be a commutative ring, and let $u_{j,{\boldsymbol{\alpha}}}\in\mathbf{k}$ for all $j=1,\ldots,r$ and $|{\boldsymbol{\alpha}}|=d_{j}$. For all $T\in A_{\mathbf{Z}}{}[\mathbf{X}]$, we denote by $T(u_{j,{\boldsymbol{\alpha}}})\in\mathbf{k}[\mathbf{X}]$ the polynomial obtained by evaluating the indeterminates $U_{j,{\boldsymbol{\alpha}}}$ at $u_{j,{\boldsymbol{\alpha}}}\in\mathbf{k}$ for all $j=1,\ldots,r$ and $|{\boldsymbol{\alpha}}|=d_{j}$ in the polynomial $T$.
(1.3) Definition.
An inertia form for the polynomials $F_{1},\ldots,F_{r}$ is an element $T\in A_{\mathbf{Z}}{}[\mathbf{X}]$ such for all $i=0,\ldots,n$, there exists $N_{i}\in\mathbf{N}$ such that $X_{i}^{N_{i}}T\in(F_{1},\ldots,F_{r})$.
In other words, the inertia forms for $F_{1},\ldots,F_{r}$ are the elements of the saturation of the ideal $(F_{1},\ldots,F_{r})$ in $A_{\mathbf{Z}}{}[\mathbf{X}]$. We denote by $\mathfrak{I}_{\mathbf{Z}}$ the ideal of all degree 0 inertia forms for $F_{1},\ldots,F_{r}$; thus, in the notation introduced in (1.1.1),
$\mathfrak{I}_{\mathbf{Z}}=\bigl(I:(X_{0},\ldots,X_{n})^{\infty}\bigr)\cap A_{\mathbf{Z}}.$
Note that this is a homogeneous ideal with respect to the standard grading on $A_{\mathbf{Z}}$. In view of the above considerations, we have the following.
(1.4) Theorem.
Suppose that $\mathbf{k}$ is a field, and let $u_{j,{\boldsymbol{\alpha}}}\in\mathbf{k}$ for for all $j=1,\ldots,r$ and $|{\boldsymbol{\alpha}}|=d_{j}$. Consider the polynomials $F_{1}(u_{1,{\boldsymbol{\alpha}}}),\ldots,F_{r}(u_{r,{\boldsymbol{\alpha}}})\in\mathbf{k}[\mathbf{X}]$. The two following propositions are equivalent:
  1. (i)
    the ideal $\bigl(F_{1}(u_{1,{\boldsymbol{\alpha}}}),\ldots,F_{r}(u_{r,{\boldsymbol{\alpha}}})\bigr)$ defines a non-empty subscheme of $\mathbf{P}^{n}_{\mathbf{k}}$;
  2. (ii)
    for all $T\in\mathfrak{I}_{\mathbf{Z}}$, $T(u_{j,{\boldsymbol{\alpha}}})=0$.
This tells us that a given specialization to a field of the polynomials $F_{j}$ defines a non-empty subscheme if and only if all the constants in the saturation of $(F_{1},\ldots,F_{r})$ vanish in this specialization (see also [6]). Note that for all $T\in\mathfrak{I}_{\mathbf{Z}}$, one has $T(u_{j,{\boldsymbol{\alpha}}})\in\mathbf{k}$, since $T$ has degree 0 in $\mathbf{X}$.
(1.5) Example.
Set $n=1$, and let $F_{1}=U_{1,0}X_{0}+U_{1,1}X_{1}$ and $F_{2}=U_{2,0}X_{0}+U_{2,1}X_{1}$. Then $R=U_{1,0}U_{2,1}-U_{1,1}U_{2,0}$ is a degree 0 inertia form of $F_{1}$ and $F_{2}$, as indeed
$X_{0}R=U_{2,1}F_{1}-U_{1,1}F_{2}\quad\text{and}\quad X_{1}R=U_{1,0}F_{2}-U_{2,0}F_{1}.$
We thus recover the usual determinant of two linear forms. We recommend that the reader come back to this example when we introduce various gradings on $A_{\mathbf{Z}}$ later on.
We emphasize that in general the subscheme of $\mathrm{Spec}(\mathbf{k})$ defined by the specialization of $\mathfrak{I}_{\mathbf{Z}}$ coincides only set-theoretically with the scheme-theoretic image of $V\to\mathrm{Spec}(\mathbf{k})$ (see [8, §3, Remarque 1] and [15, §1]), which is the reason why we assume that $\mathbf{k}$ is a field in Theorem (1.4). If $\mathbf{k}$ is an arbitrary commutative ring, what is indeed true is that the subscheme defined by the $F_{j}(u_{j,\alpha})$’s surjects onto $\mathrm{Spec}(\mathbf{k})$ as a set if and only if $\mathfrak{I}_{\mathbf{Z}}\otimes_{A_{\mathbf{Z}}}\mathbf{k}$} is contained in the nilradical $\sqrt{(0)}$ of $\mathbf{k}$, but this says nothing more than Theorem (1.4).
(1.6) Theorem (see [5, §2.1]).
The ideal of inertia forms for $F_{1},\ldots,F_{r}$ is prime, and so is the ideal $\mathfrak{I}_{\mathbf{Z}}\subseteq A_{\mathbf{Z}}$ of inertia forms of degree 0.
The resultant situation is when we consider $n+1$ homogeneous equations in $\mathbf{P}^{n}$.
(1.7) Theorem.
If $r=n+1$, the ideal $\mathfrak{I}_{\mathbf{Z}}$ is principal. Up to sign it has a unique generator, which is an irreducible element of $A_{\mathbf{Z}}$; we denote it by $\mathrm{Res}_{d_{1},\ldots,d_{n+1}}\in A_{\mathbf{Z}}$.
Moreover, for all $k\in[\kern-1.72218pt[1,n+1]\kern-1.72218pt]$, $\mathrm{Res}_{d_{1},\ldots,d_{n+1}}$ is homogeneous of degree $\prod_{j\neq k}d_{j}$ with respect to the coefficients of the polynomial $F_{k}$, i.e., with respect to the indeterminates $U_{k,{\boldsymbol{\alpha}}}$, $|{\boldsymbol{\alpha}}|=d_{k}$ (all assumed to have weight one).
Let $f_{1},\ldots,f_{n+1}\in\mathbf{k}[\mathbf{X}]$ be polynomials of respective degrees $d_{1},\ldots,d_{n+1}$. They are specializations $F_{1}(u_{1,{\boldsymbol{\alpha}}}),\ldots,F_{r}(u_{r,{\boldsymbol{\alpha}}})$ of $F_{1},\ldots,F_{n+1}\in A_{\mathbf{Z}}{}[\mathbf{X}]$ respectively, for an appropriate canonical choice of $u_{j,{\boldsymbol{\alpha}}}\in\mathbf{k}$, for all $j=1,\ldots,n+1$ and $|{\boldsymbol{\alpha}}|=d_{k}$. We let $\mathrm{Res}(f_{1},\ldots,f_{n+1})\in\mathbf{k}$ (or $\mathrm{Res}_{d_{1},\ldots,d_{n+1}}(f_{1},\ldots,f_{n+1})\in\mathbf{k}$, if we want to emphasize the dependency on the degrees) be the corresponding specialization of $\mathrm{Res}_{d_{1},\ldots,d_{n+1}}\in A_{\mathbf{Z}}$. The multi-homogeneity property stated in the above theorem may then be rephrased as follows: for all $\lambda\in\mathbf{k}$,
$\mathrm{Res}_{d_{1},\ldots,d_{n+1}}(f_{1},\ldots,\lambda f_{k},\ldots,f_{n+1})=\lambda^{d_{1}\cdots d_{k-1}d_{k+1}\cdots d_{n+1}}\,\mathrm{Res}_{d_{1},\ldots,d_{n+1}}(f_{1},\ldots,f_{k},\ldots,f_{n+1}).$
The sign indeterminacy in the definition of $\mathrm{Res}_{d_{1},\ldots,d_{n+1}}$ is usually removed by imposing the normalizing equality $\mathrm{Res}(X_{0}^{d_{0}},\ldots,X_{n}^{d_{n}})=1$.
(1.8)Divisibility property (see, e.g., [15, §5.6])
Let $F_{1},\ldots,F_{n+1}$ and $G_{1},\ldots,G_{n+1}$ be two sequences of homogeneous polynomials in $\mathbf{k}[\mathbf{X}]$ such that we have the inclusion of ideals of $\mathbf{k}[\mathbf{X}]$,
$(G_{1},\ldots,G_{n+1})\subseteq(F_{1},\ldots,F_{n+1}).$
Then, $\mathrm{Res}(F_{1},\ldots,F_{n+1})$ divides $\mathrm{Res}(G_{1},\ldots,G_{n+1})$ in $\mathbf{k}$.
Besides its ordinary multi-homogeneity property given in Theorem (1.7), the resultant has other homogeneous structures that we call “weight properties” to emphasize that the grading of the coefficient ring $A_{\mathbf{Z}}$ is not the standard one.
(1.9) Proposition (see [15, §5.13.2]).
In the notation of Paragraph (1.2), let $k$ be an integer in $\kern-1.72218pt[0,n]\kern-1.72218pt]$, and consider the grading on $A_{\mathbf{Z}}=\mathbf{Z}[U_{j,{\boldsymbol{\alpha}}}]_{j=1,\ldots,n+1,\,|{\boldsymbol{\alpha}}|=d_{j}}$ defined by
(1.9.1) $\mathrm{weight}(U_{j,{\boldsymbol{\alpha}}})=\alpha_{k}.$
In this grading, the resultant $\mathrm{Res}_{d_{1},\ldots,d_{n+1}}$ is homogeneous of degree $d_{1}d_{2}\cdots d_{n+1}$.
Bezout Theorem, which counts the number of roots of a finite complete intersection scheme in a projective space, and thus is the mother of all statements in projective enumerative geometry, can be deduced from this property. This goes as follows.
Consider $n$ homogeneous polynomials $F_{1},\ldots,F_{n}\in\mathbf{k}[X_{0},\ldots,X_{n}]$, of degrees $d_{1},\ldots,d_{n}$ respectively. The idea is to project $\mathbf{P}^{n}$ to $\mathbf{P}^{1}$ from a codimension $2$ linear space, in order to reduce to plain polynomials in one indeterminate only, i.e., homogeneous polynomials in two indeterminates. We thus consider
(1.10.1) $\tilde{F}_{j}\,(T_{0},T_{1},\ldots,T_{n-1})=F_{j}\,\bigl(T_{0}X_{0},\,T_{0}X_{1},\,T_{1},\ldots,T_{n-1}\bigr)\in\mathbf{k}[X_{0},X_{1}][T_{0},\ldots,T_{n-1}]$
for all $j=1,\ldots,n$. A point $(x_{0}:x_{1})\in\mathbf{P}^{1}_{\mathbf{k}}$ sits in the projection of $\mathrm{V}(F_{1},\ldots,F_{n})\subseteq\mathbf{P}^{n}$ if and only if the specializations of $\tilde{F}_{1},\ldots,\tilde{F}_{n}$, obtained by specializing $X_{0}$ and $X_{1}$ to $x_{0}$ and $x_{1}$ respectively, have a common zero in $\mathbf{P}^{n-1}$. We thus consider the resultant of $\tilde{F}_{1},\ldots,\tilde{F}_{n}$ with respect to the indeterminates $T_{0},\ldots,T_{n-1}$,
$\mathrm{Res}\bigl(\tilde{F}_{1},\ldots,\tilde{F}_{n}\bigr)\in\mathbf{k}[X_{0},X_{1}].$
Writing $F_{j}=\sum_{|{\boldsymbol{\alpha}}|=d_{j}}u_{j,{\boldsymbol{\alpha}}}\mathbf{X}^{\boldsymbol{\alpha}}$, one has
$\tilde{F}_{j}=\sum_{|{\boldsymbol{\alpha}}|=d_{j}}\bigl(u_{j,{\boldsymbol{\alpha}}}X_{0}^{\alpha_{0}}X_{1}^{\alpha_{1}}\bigr)\,T_{0}^{\alpha_{0}+\alpha_{1}}T_{1}^{\alpha_{2}}\cdots T_{n-1}^{\alpha_{n}}.$
Thus, $\tilde{F}_{j}$ has degree $d_{j}$ in $T_{0},\ldots,T_{n-1}$, and its coefficient in the monomial $T_{0}^{\tilde{\alpha}_{0}}T_{1}^{\tilde{\alpha}_{1}}\cdots T_{n-1}^{\tilde{\alpha}_{n-1}}$ is a degree $\tilde{\alpha}_{0}$ homogeneous polynomial in $X_{0},X_{1}$.
Then, it follows from the above Proposition (1.9) that
$\mathrm{Res}_{d_{1},\ldots,d_{n}}\bigl(\tilde{F}_{1},\ldots,\tilde{F}_{n}\bigr)\in\mathbf{k}[X_{0},X_{1}]$
is homogeneous of degree $d_{1}\cdots d_{n}$. Therefore, $\mathrm{V}\bigl(\mathrm{Res}(\tilde{F}_{1},\ldots,\tilde{F}_{n})\bigr)\subseteq\mathbf{P}^{1}_{\mathbf{k}}$ consists of $d_{1}\cdots d_{n}$ points, counted with multiplicities. One may then conclude that $\mathrm{V}(F_{1},\ldots,F_{n})\subseteq\mathbf{P}^{n}_{\mathbf{k}}$ itself consists of $d_{1}\cdots d_{n}$ points, by the classical arguments used to prove Bezout’s theorem for the intersection of two plane curves, see e.g., [2, §4.5].
The following is a parent of the weight property given in Proposition (1.9) above. It is proven in [15, §5.13] as well.
(1.11) Proposition.
In the same situation as in Proposition (1.9) above, consider the grading on $A_{\mathbf{Z}}$ defined by
($1.1$) $\mathrm{weight}(U_{j,{\boldsymbol{\alpha}}})=d_{j}-\alpha_{k}.$
In this grading, the resultant $\mathrm{Res}_{d_{1},\ldots,d_{n+1}}\in A_{\mathbf{Z}}$ is homogeneous of degree $nd_{1}d_{2}\cdots d_{n+1}$.
One may also use the following corollary, based on the standard homogeneity of the resultant in Theorem (1.7), to obtain additional weight properties from (1.9) and (1.11).
(1.12) Corollary.
Assume that the resultant $\mathrm{Res}_{d_{1},\ldots,d_{n+1}}$ is homogeneous of degree $\delta$ for the grading on $A_{\mathbf{Z}}$ defined by $\mathrm{weight}(U_{j,{\boldsymbol{\alpha}}})=w_{j,{\boldsymbol{\alpha}}}$. Let $r_{1},\ldots,r_{n+1}\in\mathbf{Z}$. For the new grading on $A_{\mathbf{Z}}$ defined by $\mathrm{weight}(U_{j,{\boldsymbol{\alpha}}})=w_{j,{\boldsymbol{\alpha}}}+r_{j}$, the resultant $\mathrm{Res}_{d_{1},\ldots,d_{n+1}}$ is homogeneous of degree
$\delta+\sum_{1\leqslant k\leqslant n+1}\Bigl(r_{k}\,\prod\nolimits_{j\neq k}d_{j}\Bigr).$
Proof.
Let $k\in[\kern-1.72218pt[1,n+1]\kern-1.72218pt]$. Since the resultant is homogeneous of degree $\prod_{j\neq k}d_{j}$ with respect to the indeterminates $(U_{k,{\boldsymbol{\alpha}}})_{|{\boldsymbol{\alpha}}|=d_{k}}$ (for the standard grading), a shift by $r_{k}$ in the weights of all the indeterminates $(U_{k,{\boldsymbol{\alpha}}})_{|{\boldsymbol{\alpha}}|=d_{k}}$ induces a shift by $r_{k}\,\prod_{j\neq k}d_{j}$ in the degree of the resultant. ∎
The reduced resultant
We shall now explain how to adapt the ideas of the previous paragraph to develop the theory of the reduced resultant. We refer to [21] and [16] for the details and proofs. Somehow, this is a generalization of the following toy example.
(1.1) Example (projection of a complete intersection from one of its points).
Let $F,G\in\mathbf{k}[\mathbf{X}]$ be two homogeneous polynomials of degrees $a$ and $b$, defining a complete intersection $V\subseteq\mathbf{P}^{n}$, and suppose one wants to project $V$ from a point $p_{0}\in\mathbf{P}^{n}$. Assume for simplicity that $\mathbf{k}$ is an algebraically closed field. We may take $p_{0}=(1:0:\ldots:0)$. Then, we consider the two polynomials
(1.1.1) $F(T,X_{1},\ldots,X_{n})$ $=F_{0}T^{a}+F_{1}T^{a-1}+\cdots+F_{a}$
$\text{and}\quad G(T,X_{1},\ldots,X_{n})$ $=G_{0}T^{b}+G_{1}T^{b-1}+\cdots+G_{b}$
in $\mathbf{k}[X_{1},\ldots,X_{n}][T]$. We are abusing notation here, as one should consider instead the two polynomials $F(T,S\,X_{1},\ldots,S\,X_{n})$ and $G(T,S\,X_{1},\ldots,S\,X_{n})$ that are homogeneous in the couple of indeterminates $(S,T)$. Moreover, beware that (1.1.1) above is written “in reverse order” with respect to what we did when discussing Bezout Theorem, see (1.10.1).
If $p_{0}\not\in V$, the point $(x_{1}:\ldots:x_{n})\in\mathbf{P}_{\mathbf{k}}^{n-1}$ belongs to the projection of $V$ from $p_{0}$ if and only if the two polynomials in (1.1.1) have a common root in $\mathbf{P}^{1}$, hence the equation of the projection is given by
$\mathrm{Res}_{a,b}(F,G)\in\mathbf{k}[X_{1},\ldots,X_{n}],$
which is homogeneous of degree $ab$ in the indeterminates $(X_{1},\ldots,X_{n})$ by Proposition (1.11), as the polynomials $F_{i}$ are homogeneous of degree $i$ in $\mathbf{k}[X_{1},\ldots,X_{n}]$.
On the other hand, if $p_{0}\in V$ then, letting $a^{\prime}$ and $b^{\prime}$ be the respective multiplicities of $p_{0}$ in the hypersurfaces $\mathrm{V}(F)$ and $\mathrm{V}(G)$, one has
$F_{0}=\cdots=F_{a^{\prime}-1}=G_{0}=\cdots=G_{b^{\prime}-1}=0,$
so that (1.1.1) becomes
(1.1.2) $F(T,X_{1},\ldots,X_{n})$ $=F_{a^{\prime}}T^{a-a^{\prime}}+\cdots+F_{a}=\vphantom{F}^{\flat}\kern-1.00006ptF$
$\text{and}\quad G(T,X_{1},\ldots,X_{n})$ $=G_{b^{\prime}}T^{b-b^{\prime}}+\cdots+G_{b}=\vphantom{G}^{\flat}\kern-1.00006ptG.$
It follows that the equation of the projection of $V$ from $p_{0}$ is given by
$\mathrm{Res}_{a-a^{\prime},b-b^{\prime}}(\vphantom{F}^{\flat}\kern-1.00006ptF,\vphantom{G}^{\flat}\kern-1.00006ptG)\in\mathbf{k}[X_{1},\ldots,X_{n}].$
We shall see later on that this polynomial is the reduced resultant of $F$ and $G$ truncated at orders $a-a^{\prime}$ and $b-b^{\prime}$ respectively, as polynomials in the indeterminate $T$. It is a homogeneous polynomial of degree $ab-a^{\prime}b^{\prime}$ in $(X_{1},\ldots,X_{n})$: indeed, the coefficient of $\vphantom{F}{}^{\flat}\kern-1.00006ptF$ (respectively, $\vphantom{G}{}^{\flat}\kern-1.00006ptG$) in $T^{i}$ is a homogeneous polynomial in $(X_{1},\ldots,X_{n})$ of degree $a-i=\deg_{T}(\vphantom{F}^{\flat}\kern-1.00006ptF)-i+a^{\prime}$ (respectively, $b-i=\deg_{T}(\vphantom{G}^{\flat}\kern-1.00006ptG)-i+b^{\prime}$). Therefore, Corollary (1.12) applied to the grading of Proposition (1.11) gives that $\mathrm{Res}_{a-a^{\prime},b-b^{\prime}}(\vphantom{F}^{\flat}\kern-1.00006ptF,\vphantom{G}^{\flat}\kern-1.00006ptG)$ is homogeneous of degree
$(a-a^{\prime})(b-b^{\prime})+a^{\prime}(b-b^{\prime})+b^{\prime}(a-a^{\prime})=ab-a^{\prime}b^{\prime},$
as we had announced. This weight property is an instance of that given in Corollary (1.4), which applies to reduced resultants in general.
Let $d_{1},\ldots,d_{n+1}$ be positive integers. For all $j\in[\kern-1.72218pt[1,n+1]\kern-1.72218pt]$, we consider the generic homogeneous degree $d_{j}$ polynomial $F_{j}$, which we write as
$F_{j}=\sum_{|{\boldsymbol{\alpha}}|=d_{j}}U_{j,{\boldsymbol{\alpha}}}\mathbf{X}^{\boldsymbol{\alpha}}=\sum_{k=0}^{d_{j}}X_{0}^{d_{j}-k}\,F_{j,k}(X_{1},\ldots,X_{n}).$
Thus, for all $k=0,\ldots,d_{j}$, the polynomial $F_{j,k}$ is a degree $k$ homogeneous element of $A_{\mathbf{Z}}{}[X_{1},\ldots,X_{n}]$. We fix integers $s_{j}\in[\kern-1.72218pt[1,d_{j}]\kern-1.72218pt]$ for all $j=1,\ldots,n+1$. The truncation of $F_{j}$ at order $d_{j}-s_{j}$ with respect to $X_{0}$ is the polynomial
$H_{j}=\sum_{k=s_{j}}^{d_{j}}X_{0}^{d_{j}-k}F_{j,k}=X_{0}^{d_{j}-s_{j}}F_{j,s_{j}}+\cdots+X_{0}F_{j,d_{j}-1}+F_{j,d_{j}}\in A_{\mathbf{Z}}{}[\mathbf{x}].$
This definition gives a special role to the indeterminate $X_{0}$, and to the point $(1:0:...:0)\in\mathbf{P}^{n}$. One may thus think of the latter as the (chosen) origin in the affine space $\mathbf{P}^{n}-\mathrm{V}(X_{0})$, and of $\mathrm{V}(X_{0})$ as the hyperplane at infinity. We may occasionally use this terminology.
The purpose of reduced elimination theory is to study inertia forms of the truncations $H_{1},\ldots,H_{n+1}$ defined above. Note that, for all $j=1,\ldots,n+1$, the truncation $H_{j}$ is the generic homogeneous polynomial of degree $d_{j}$ with a multiplicity $s_{j}$ zero in the origin (1:0::0)(1:0:...:0). The wish for the reduced resultant is that it is a polynomial in the coefficients of $H_{1},\ldots,H_{n+1}$, which vanishes if and only if $H_{1},\ldots,H_{n+1}$ have a non-trivial common root in addition to that of multiplicity $s_{1}\cdots s_{n+1}$ at the origin.
This can be done with essentially the same strategy as in the classical case, which we have reviewed in Section 1.1. We will only consider the “resultant situation”, when the number of polynomials is $n+1$.
(1.3) Theorem ([21, Theorem 6 and §8] and [16, Theorem II.0.5 and §IV.0]).
Assume that $d_{j}>s_{j}$ for some $j\in[\kern-1.72218pt[1,n+1]\kern-1.72218pt]$. The ideal of reduced inertia forms of degree 0
${\mathcal{Q}}_{\mathbf{Z}}=\bigl((H_{1},\ldots,H_{n+1}):(X_{1},\ldots,X_{n})^{\infty}\bigr)\cap A_{\mathbf{Z}}$
is a prime and principal ideal of $A_{\mathbf{Z}}{}$. The reduced resultant, denoted
$\operatorname{redRes}_{d_{1},\ldots,d_{n+1}}^{s_{1},\ldots,s_{n+1}}\in A_{\mathbf{Z}},$
is defined, up to sign, as the generator of ${\mathcal{Q}}_{\mathbf{Z}}$; it is therefore an irreducible element of $A_{\mathbf{Z}}{}$.
Moreover, if $d_{j}>s_{j}$ for at least two distinct integers $j,j^{\prime}\in[\kern-1.72218pt[1,n+1]\kern-1.72218pt]$, then for all $i\in[\kern-1.72218pt[1,n+1]\kern-1.72218pt]$ the reduced resultant is a homogeneous polynomial of degree
$\frac{d_{1}d_{2}\cdots d_{n+1}}{d_{i}}-\frac{s_{1}s_{2}\cdots s_{n+1}}{s_{i}}$
with respect to the coefficients of the polynomial $H_{i}$, i.e., with respect to the indeterminates $U_{i,{\boldsymbol{\alpha}}}$ such that $|{\boldsymbol{\alpha}}|=d_{i}$ and $\alpha_{0}\leqslant d_{i}-s_{i}$.
If there is only one integer $j\in[\kern-1.72218pt[1,n+1]\kern-1.72218pt]$ such that $d_{j}>s_{j}$, then the reduced resultant is equal to the resultant of the polynomials $H_{1},\ldots,H_{j-1},\,H_{j+1},\ldots,H_{n+1}$, as polynomials in $X_{1},\ldots,X_{n}$ only.222if $s_{j}=d_{j}$, then $H_{j}=F_{j,d_{j}}$ is a polynomial in $X_{1},\ldots,X_{n}$ only; geometrically, the polynomial $H_{j}$ defines a cone in $\mathbf{P}^{n}$ over the hypersurface defined by $F_{j,d_{j}}$ in $\mathbf{P}^{n-1}$.
The sign indeterminacy in the definition of the reduced resultant can be removed by means of Theorem (1.3) below, once the sign of the ordinary resultant has been chosen.
Note that, in the above statement, the ideal $(H_{1},\ldots,H_{n+1})$ is saturated with respect to $(X_{1},\ldots,X_{n})$, which is the defining ideal of the point $(1:0:\ldots:0)$, whereas for the plain resultant we considered instead the saturation with respect to the irrelevant ideal $(X_{0},X_{1},\ldots,X_{n})$. Beware moreover that the polynomials $H_{1},\ldots,H_{n+1}$ are not homogeneous in the set of indeterminates $(X_{1},\ldots,X_{n})$.
The reduced resultant depends, of course, only on the coefficients of the generic truncated polynomials $H_{1},\ldots,H_{n+1}$, and not on all the coefficients of the polynomials $F_{1},\ldots,F_{n+1}$. We will often denote it by $\operatorname{redRes}(H_{1},\ldots,H_{n+1})$ without printing the integers $d_{i}$ and $s_{i}$, that are implicitly given by the polynomials $H_{1},\ldots,H_{n+1}$. We may also use the notation $\operatorname{redRes}(F_{1},\ldots,F_{n+1})$, to avoid giving particular names to the truncation.
The reduced resultant of a collection of polynomials $h_{1},\ldots,h_{n+1}\in\mathbf{k}[\mathbf{X}]$ with zeros of respective multiplicities at least $s_{1},\ldots,s_{n+1}$ at the origin is defined as the corresponding specialization of the generic reduced resultant; it is an element in $\mathbf{k}$, denoted by $\operatorname{redRes}(h_{1},\ldots,h_{n+1})$, or possibly $\operatorname{redRes}(f_{1},\ldots,f_{n+1})$ depending on the context.
(1.4)Vanishing of the reduced resultant
The reduced resultant $\operatorname{redRes}(H_{1},\ldots,H_{n+1})$ is a polynomial in the coefficients of the polynomials $H_{j}$, $j=1,\ldots,n+1$, i.e., an element of the ring
$\mathbf{Z}\bigl[U_{i,{\boldsymbol{\alpha}}}\bigr]_{i=1,\ldots,n+1,\,|{\boldsymbol{\alpha}}|=d_{i},\,\alpha_{0}\leqslant d_{i}-s_{i}}\subseteq A_{\mathbf{Z}}{}.$
Its vanishing on an algebraically closed field $\mathbf{k}$ characterizes those collections of hypersurfaces of $\mathbf{P}^{n}_{\mathbf{k}}$ defined by $h_{1},\ldots,h_{n+1}$ that have a further intersection point, infinitely near or not, besides the origin $p_{0}=(1\!:\!0\!:\!\ldots\!:\!0)$, i.e., those collections such that one of the two following conditions holds:
  1. (a)
    the hypersurfaces defined by $h_{1},\ldots,h_{n+1}$ intersect at a point which is different from $p_{0}$;
  2. (b)
    the polynomials $f_{j,s_{j}}$, $j=1,\ldots,n+1$, have a common root in $\mathbf{P}^{n-1}_{\mathbf{k}}$, which means that the tangent cones of the hypersurfaces $\mathrm{V}(h_{1}),\ldots,\mathrm{V}(h_{n+1})$ at $p_{0}$ have a line in common; equivalently, the proper transforms of $\mathrm{V}(h_{1}),\ldots,\mathrm{V}(h_{n+1})$ in the blow-up of $\mathbf{P}^{n}_{\mathbf{k}}$ at $p_{0}$ have a point in common which lies on the exceptional divisor.
This property of the reduced resultant is proved in [21, Theorem 3.1, 3.2 and 3.3], and in [16, Proposition I.1].
Generalized weight properties
In [21] Zariski showed that the reduced resultant can be computed from the corresponding resultant. To obtain this property, he introduced a generalization of the grading (1.9.1) and, although the resultant is not homogeneous with respect to this new grading, he proved that its graded piece of smallest degree is connected to the reduced resultant.
We maintain the notation of Section 1.2.
(1.1)The Zariski grading
We define a grading on $A_{\mathbf{Z}}=\mathbf{Z}[U_{j,{\boldsymbol{\alpha}}}]$ by assigning for all $j$
$\mathrm{weight}(U_{j,{\boldsymbol{\alpha}}})=\begin{cases}0&\text{if}\ \alpha_{0}<d_{j}-s_{j}\\ \alpha_{0}-d_{j}+s_{j}&\text{otherwise},\end{cases}$
and weight 0 to the constants. We find it helpful to visualize this definition as follows:
[…]
Note in particular that the indeterminates $U_{j,{\boldsymbol{\alpha}}}$ whose weight equals 0 in this grading are exactly the coefficients of the truncation $H_{j}$ of the polynomial $F_{j}$. The grading (1.9.1) introduced in Proposition (1.9) is a particular case of a Zariski grading (corresponding to $s_{j}=d_{j}$ for all $j$), which explains the terminology “generalized weight properties”.
The main property of the Zariski grading is that it allows the computation of the reduced resultant of $H_{1},\ldots,H_{n+1}$ (the truncations of $F_{1},\ldots,F_{n+1}$ at the orders $d_{1}-s_{1},\ldots,d_{n+1}-s_{n+1}$, respectively) from the resultant of $F_{1},\ldots,F_{n+1}$. We need one more piece of notation to see how this goes.
For all $j=1,\ldots,n+1$, we let $G_{j}$ be the quotient of the Euclidean division of $F_{j}$ by $X_{0}^{d_{j}-s_{j}}$ in $A_{\mathbf{Z}}{}[X_{1},\ldots,X_{n}][X_{0}]$, i.e.,
$G_{j}\,=\,\frac{1}{X_{0}^{d_{j}-s_{j}}}\sum_{k\leqslant s_{j}}X_{0}^{d_{j}-k}\,F_{j,k}\,=\,\sum_{k=0}^{s_{j}}X_{0}^{s_{j}-k}\,F_{j,k}\\ =X_{0}^{s_{j}}\,F_{j,0}+\cdots+X_{0}\,F_{j,s_{j}-1}+F_{j,s_{j}}.$
The polynomial $G_{j}$ is a generic degree $s_{j}$ homogeneous polynomial in the set of variables $\mathbf{X}$. Beware that $F_{j}$ does not equal $H_{j}+X_{0}^{d_{j}-s_{j}}G_{j}$, as $F_{j,s_{j}}$ appears in both $H_{j}$ and $G_{j}$:
[…]
Also, we advise the reader to compare Displays (1.1.1) and (1.2.1).
(1.3) Theorem (Zariski Formula, [21, Theorem 5.1 and Theorem 5.2] and [16, Lemme IV.1.6]).
The nonzero homogeneous piece of lowest degree with respect to the Zariski grading of $\mathrm{Res}(F_{1},\ldots,F_{n+1})$ has degree $s_{1}s_{2}\cdots s_{n+1}$. Denote it by $\bigl[\mathrm{Res}(F_{1},\ldots,F_{n+1})\bigr]_{s_{1}s_{2}\cdots s_{n+1}}$.
(a) If $s_{j}<d_{j}$ for at least two distinct integers $j=j_{1},j_{2}\in[\kern-1.72218pt[1,n+1]\kern-1.72218pt]$, then
$\bigl[\mathrm{Res}(F_{1},\ldots,F_{n+1})\bigr]_{s_{1}s_{2}\cdots s_{n+1}}=\mathrm{Res}(G_{1},\ldots,G_{n+1})\cdot\operatorname{redRes}(H_{1},\ldots,H_{n+1}).$
(b) If there exists $j_{0}\in[\kern-1.72218pt[1,n+1]\kern-1.72218pt]$ such that $s_{j_{0}}<d_{j_{0}}$ and $s_{j}=d_{j}$ for all $j\neq j_{0}$, then
$\bigl[\mathrm{Res}(F_{1},\ldots,F_{n+1})\bigr]_{s_{1}s_{2}\cdots s_{n+1}}=\mathrm{Res}(G_{1},\ldots,G_{n+1})\cdot\operatorname{redRes}(H_{1},\ldots,H_{n+1})^{d_{j_{0}}-s_{j_{0}}}.$
Note that despite appearances, there is a Zariski Formula in all cases, because if none of the two conditions in (a) and (b) of the above statement is verified, then the reduced resultant is actually an ordinary resultant.
Observe that, in the Zariski grading: (i) the coefficients of $H_{1},\ldots,H_{n+1}$ all have weight 0, hence $\operatorname{redRes}(H_{1},\ldots,H_{n+1})$ is homogeneous of degree 0; (ii) for all $j=1,\ldots,n+1$, the coefficient of $G_{j}$ in $\mathbf{X}^{\boldsymbol{\alpha}}$ has weight $\alpha_{0}$ hence, by Proposition (1.9), $\mathrm{Res}(G_{1},\ldots,G_{n+1})$ is homogeneous of weight $s_{1}s_{2}\cdots s_{n+1}$. Thus, indeed,
$\mathrm{Res}(G_{1},\ldots,G_{n+1})\cdot\operatorname{redRes}(H_{1},\ldots,H_{n+1})^{e}$
is homogeneous of weight $s_{1}s_{2}\cdots s_{n+1}$ for all $e$.
(1.4) Corollary.
Consider the grading of $A_{\mathbf{Z}}$ defined by
$\mathrm{weight}(U_{j,{\boldsymbol{\alpha}}})=d_{j}-\alpha_{0}.$
With respect to this grading, $\operatorname{redRes}(H_{1},\ldots,H_{n+1})$ is homogeneous of degree
$n(d_{1}\cdots d_{n+1}-s_{1}\cdots s_{n+1})$
if there are at least two integers $j=j_{1},j_{2}$ such that $s_{j}<d_{j}$, and homogeneous of degree
$n\,d_{1}\ldots d_{n+1}/d_{j_{0}}$
if there exists $j_{0}$ such that $s_{j_{0}}<d_{j_{0}}$ and $s_{j}=d_{j}$ for all $j\neq j_{0}$.
This is the homogeneity property of the reduced resultant that particularizes to give the degree of $\mathrm{Res}(\vphantom{F}^{\flat}\kern-1.00006ptF,\vphantom{G}^{\flat}\kern-1.00006ptG)$ in Example (1.1).
Proof.
The grading of $A_{\mathbf{Z}}$ we are considering is that of Proposition (1.11), with $k=0$; thus for all $j=1,\ldots,n+1$ the coefficients of $F_{j,l}$ all have weight $l$, for $l=0,\ldots,d_{j}$. Then we know by Proposition (1.11) that the resultant of $F_{1},\ldots,F_{n+1}$ is homogeneous of degree $n\,d_{1}\cdots d_{n+1}$, and by the same argument the resultant of $G_{1},\ldots,G_{n+1}$ is homogeneous of degree $n\,s_{1}\cdots s_{n+1}$. Therefore, we deduce from Theorem (1.3) that the reduced resultant of $H_{1},\ldots,H_{n+1}$ is homogeneous, of degree
$n\,d_{1}\cdots d_{n+1}-n\,s_{1}\cdots s_{n+1}$
if $s_{j}<d_{j}$ for two distinct indices $j=j_{1},j_{2}$, and of degree
$\frac{n\,d_{1}\cdots d_{n+1}-n\,s_{1}\cdots s_{n+1}}{d_{j_{0}}-s_{j_{0}}}=n\,\frac{d_{1}\cdots d_{n+1}}{d_{j_{0}}}\,\frac{d_{j_{0}}-s_{j_{0}}}{d_{j_{0}}-s_{j_{0}}}$
if $s_{j}=d_{j}$ for all $j\neq j_{0}$ and $s_{j_{0}}<d_{j_{0}}$. ∎
Reduced discriminant and Salmon formula
In this Section we give a rigorous proof of formula (a). This is done by introducing the concept of reduced discriminant. We begin with a quick recap on the ordinary discriminant of a hypersurface, following [5, §4]; see also [8] and [12, Chapter 13, §D].
Discriminant of a homogeneous polynomial
Let $d$ be a positive integer, and consider the generic homogeneous degree $d$ polynomial $F=\sum_{|{\boldsymbol{\alpha}}|=d}U_{\alpha}\mathbf{X}^{\boldsymbol{\alpha}}$ in $n+1$ indeterminates $\mathbf{X}=(X_{0},\ldots,X_{n})$. We set $A_{\mathbf{Z}}=\mathbf{Z}[U_{{\boldsymbol{\alpha}}}]_{|{\boldsymbol{\alpha}}|=d}$. For all $i=0,\ldots,n$ we let $\partial_{i}$ denote derivation with respect to the indeterminate $X_{i}$.
(2.2) Definition.
There is a unique element $\mathrm{Disc}_{d}(F)\in A_{\mathbf{Z}}$ (often simply denoted by $\mathrm{Disc}(F)$) such that
(2.2.1) $d^{a(n,d)}\mathrm{Disc}_{d}(F)=\mathrm{Res}\bigl(\partial_{0}F,\ldots,\partial_{n}F\bigr)$
in $A_{\mathbf{Z}}{}$, where $a(n,d)=\frac{(d-1)^{n+1}-(-1)^{n+1}}{d}\in\mathbf{Z}$. It is homogeneous of degree $(n+1)(d-1)^{n}$ with respect to the coefficients of the polynomial $F$, i.e., with respect to the indeterminates $U_{{\boldsymbol{\alpha}}}$, $|{\boldsymbol{\alpha}}|=d$.
For a homogeneous degree $d$ polynomial $f\in\mathbf{k}[\mathbf{X}]$, we define the discriminant $\mathrm{Disc}(f)\in\mathbf{k}$ of $f$ as the specialization $\sigma(\mathrm{Disc}(F))\in\mathbf{k}$, where $\sigma:A_{\mathbf{Z}}\to\mathbf{k}$ is the unique specialization morphism mapping $F$ to $f$.
We emphasize that the factor $d^{a(n,d)}$ in Equation (2.2.1) is here to make the discriminant a universal object (i.e., to make it behave well under specialization) which yields the expected smoothness criterion. An alternative way of defining the discriminant is to use Formula (2.5.1) below. Either way there is an annoying parasitic factor. The issue comes from the fact that, while the resultant characterizes the existence of a non-trivial common zero in $\mathbf{P}^{n}$ for $n+1$ polynomials, the wish for the discriminant is that it will characterize the existence of a non-trivial common zero for the $n+2$ polynomials $F,\partial_{0}F,\partial_{1}F,\ldots,\partial_{n}F$ which are linked by the Euler Formula. The next result tells us that with Definition (2.2), the discriminant grants our wish.
(2.3) Proposition.
The ideal of inertia forms
$(\partial_{0}F,\partial_{1}F,\ldots,\partial_{n}F,F):(X_{0},\ldots,X_{n})^{\infty}$
is a prime and principal ideal in $A_{\mathbf{Z}}$. It is generated by the discriminant $\mathrm{Disc}(F)$, which is therefore an irreducible polynomial in $A_{\mathbf{Z}}$.
In fact, what we really want is that the (non-)vanishing of the discriminant characterizes smoothness. By the Jacobian criterion for smoothness, this is equivalent to the non-existence of a common root $F,\partial_{0}F,\partial_{1}F,\ldots,\partial_{n}F$. Thus, Proposition (2.3) gives the following.
(2.4) Theorem (smoothness criterion).
Suppose $\mathbf{k}$ is an algebraically closed field, and consider a degree $d$ homogeneous polynomial $f\in\mathbf{k}[\mathbf{X}]$. The following are equivalent:
  1. (i)
    the hypersurface $\mathrm{V}(f)\subseteq\mathbf{P}^{n}_{\mathbf{k}}$ is smooth;
  2. (ii)
    $\mathrm{Disc}(f)\neq 0$.
The following formula gives an alternative way of defining the discriminant. This is the path we shall follow to define the reduced discriminant. As we shall see, however, there will be some additional technical difficulty to overcome, which did not occur when passing from the ordinary resultant to its reduced form either.
(2.5) Proposition.
Let $\bar{F}$ be the polynomial $F(0,X_{1},\ldots,X_{n})\in A_{\mathbf{Z}}{}[X_{1},\ldots,X_{n}]$. We have the following identity in $A_{\mathbf{Z}}$:
(2.5.1) $\mathrm{Res}\bigl(\partial_{1}F,\ldots,\partial_{n}F,F\bigr)=\mathrm{Disc}(F)\cdot\mathrm{Disc}(\bar{F}).$
The polynomial $\bar{F}$ is the equation of the hypersurface in $\mathbf{P}^{n-1}$ cut out by $\mathrm{V}(F)$ on the hyperplane $\mathrm{V}(X_{0})\subseteq\mathbf{P}^{n}$, of which we think as the hyperplane at infinity. If $\mathbf{k}$ is an algebraically closed field and $f\in\mathbf{k}[\mathbf{X}]$ is a degree $d$ homogeneous polynomial, then, by Theorem (2.4), the vanishing of $\mathrm{Disc}(\bar{f})$ is equivalent to the hyperplane section at infinity $\mathrm{V}(f)\cap\mathrm{V}(X_{0})$ being singular. For a general $f$ such that $\mathrm{Disc}(\bar{f})=0$, the hypersurface $\mathrm{V}(f)$ is non-singular and tangent to the hyperplane $\mathrm{V}(X_{0})$.
Proof of Proposition (2.5).
We shall prove the following identity in $A_{\mathbf{Z}}$:
(2.5.2) $d^{(d-1)^{n}}\mathrm{Res}\bigl(\partial_{1}F,\ldots,\partial_{n}F,F\bigr)=\mathrm{Res}\bigl(\partial_{0}F,\ldots,\partial_{n}F\bigr)\cdot\mathrm{Res}\bigl(\partial_{1}\bar{F},\ldots,\partial_{n}\bar{F}\bigr).$
We claim that it is equivalent to Formula (2.5.1). To see why, first note that at the right-hand-side of Display (2.5.2), the first (respectively second) factor is the resultant of $n+1$ (respectively $n$) polynomials in $n+1$ (respectively $n$) indeterminates. Since $a(n,d)+a(n-1,d)=(d-1)^{n}$, it follows from Formula (2.2.1) that the identities (2.5.2) and (2.5.1) are indeed equivalent.
Now, the identity (2.5.2) may be derived from the Euler Formula, as follows. The latter formula says that $d\cdot F$ is congruent to $X_{0}\,\partial_{0}F$ modulo the ideal $(\partial_{1}F,\ldots,\partial_{n}F)$, hence
(2.5.3) $\mathrm{Res}\bigl(\partial_{1}F,\ldots,\partial_{n}F,\,d\cdot F\bigr)=\mathrm{Res}\bigl(\partial_{1}F,\ldots,\partial_{n}F,\,X_{0}\,\partial_{0}F\bigr)$
by [15, §5.9, Transformations élémentaires]. The left-hand-side of (2.5.3) equals that of (2.5.2) by the elementary homogeneity properties of the resultant, see Theorem (1.7). Its right-hand-side, on the other hand, equals
$\mathrm{Res}\bigl(\partial_{1}F,\ldots,\partial_{n}F,\,X_{0}\bigr)\cdot\mathrm{Res}\bigl(\partial_{1}F,\ldots,\partial_{n}F,\,\partial_{0}F\bigr)$
by multiplicativity of the resultant, see [15, §5.7]. Eventually,
$\mathrm{Res}\bigl(\partial_{1}F,\ldots,\partial_{n}F,\,X_{0}\bigr)=(-1)^{n(d-1)^{n}}\mathrm{Res}\bigl(\left.\partial_{1}F\right|_{X_{0}=0},\ldots,\left.\partial_{n}F\right|_{X_{0}=0}\bigr)$
by [15, (5.13.5)]333the $(-1)^{n(d-1)^{n}}$ factor is due to the fact that we need to exchange the roles of $X_{0}$ and $X_{n}$, which is not indifferent since by definition the resultant is normalized by imposing $\mathrm{Res}(X_{0}^{d_{0}},\ldots,X_{n}^{d_{n}})=1$, see right after Theorem (1.7). To work out the claimed identity, one may first perform a base change by exchanging $X_{0}$ and $X_{n}$ which, by [15, (5.13.1)], gives $\mathrm{Res}\bigl(\partial_{1}F,\ldots,\partial_{n}F,\,X_{0}\bigr)=(-1)^{(d-1)^{n}}\mathrm{Res}\bigl(\partial_{1}F(X_{n},X_{1},\ldots,X_{n-1},X_{0}),\,\ldots,\ X_{n}\bigr);$ then, apply [15, (5.13.5)] to the resultant on the right-hand-side, and eventually perform another base change to go from $(X_{1},\ldots,X_{n-1},X_{0})$ to $(X_{0},X_{1},\ldots,X_{n-1})$, which introduces the additional factor $(-1)^{(n-1)(d-1)^{n}}$, hence in total the factor $(-1)^{n(d-1)^{n}}$ indeed. An easier way to get the correct sign in general is to consider the normalizing condition, once one knows that there exists $c\in\mathbf{Z}$ such that $\mathrm{Res}_{d_{1},\ldots,d_{n},1}(F_{1},\ldots,F_{n},X_{0})=c\,\mathrm{Res}_{d_{1},\ldots,d_{n}}(\left.F_{1}\right|_{X_{0}=0},\ldots,\left.F_{n}\right|_{X_{0}=0}).$ Specializing $F_{1},\ldots,F_{n}$ to $X_{1}^{d_{1}},\ldots,X_{n}^{d_{n}}$ respectively, one gets $\mathrm{Res}_{d_{1},\ldots,d_{n},1}(X_{1}^{d_{1}},\ldots,X_{n}^{d_{n}},X_{0})=c\,\mathrm{Res}_{d_{1},\ldots,d_{n}}(X_{1}^{d_{1}},\ldots,X_{n}^{d_{n}}),$ where the left-hand-side equals $(-1)^{nd_{1}\cdots d_{n}}$ by [15, §5.8], and the right-hand-side equals $c$. , and
$\mathrm{Res}\bigl(\partial_{1}F,\ldots,\partial_{n}F,\,\partial_{0}F\bigr)=(-1)^{n(d-1)^{n+1}}\mathrm{Res}\bigl(\partial_{0}F,\,\partial_{1}F,\ldots,\partial_{n}F\bigr)$
by [15, §5.8, Effet d’une permutation des polynômes]. The upshot is that
$\mathrm{Res}\bigl(\partial_{1}F,\ldots,\partial_{n}F,\,X_{0}\,\partial_{0}F\bigr)=(-1)^{nd(d-1)^{n}}\,\mathrm{Res}\bigl(\partial_{0}F,\ldots,\partial_{n}F\bigr)\cdot\mathrm{Res}\bigl(\partial_{1}\bar{F},\ldots,\partial_{n}\bar{F}\bigr),$
which gives the result since d(d1)d(d-1) is always even. ∎
Similarly to the resultant, the discriminant is homogeneous under the two gradings of the coefficient ring $A_{\mathbf{Z}}$ introduced in Propositions (1.9) and (1.11).
(2.6) Proposition.
Let $k$ be an integer in $\kern-1.72218pt[0,n]\kern-1.72218pt]$.
In the grading of $A_{\mathbf{Z}}$ defined by $\mathrm{weight}(U_{\boldsymbol{\alpha}})=\alpha_{k}$, the discriminant $\mathrm{Disc}(F)$ is homogeneous of degree $d(d-1)^{n}$.
In the grading of $A_{\mathbf{Z}}$ defined by $\mathrm{weight}(U_{\boldsymbol{\alpha}})=d-\alpha_{k}$, the discriminant $\mathrm{Disc}(F)$ is homogeneous of degree $nd(d-1)^{n}$.
One may use (2.6.1) to compute the degree of the dual to a smooth hypersurface in $\mathbf{P}^{n+1}$, see Paragraph (3.2).
Proof.
The idea is to consider (2.2.1) and the corresponding weight properties of the resultant. We may assume $k\neq 0$, for otherwise we may write a formula analogous to (2.5.1) with respect to an indeterminate other than $X_{0}$ in order to reduce to this case.
First consider the grading of (2.6.1). For $j\neq k$ the coefficient of $\partial_{j}F$ in $\mathbf{X}^{\boldsymbol{\alpha}}$ has weight $\alpha_{k}$, whereas for $j=k$ the coefficient of $\partial_{k}F$ in $\mathbf{X}^{\boldsymbol{\alpha}}$ has weight $\alpha_{k}+1$. We may thus apply Corollary (1.12) to the homogeneity property (1.9), with $r_{k}=1$ and $r_{j}=0$ for all $j\neq k$, which gives that $\mathrm{Res}\bigl(\partial_{0}F,\ldots,\partial_{n}F\bigr)$ is homogeneous of degree
$(d-1)^{n+1}+(d-1)^{n}.$
Using (2.2.1), this gives (2.6.1).
Similarly, for the grading of (2.6.2): for $j\neq k$ the coefficient of $\partial_{j}F$ in $\mathbf{X}^{\boldsymbol{\alpha}}$ has weight $d-\alpha_{k}=(d-1)-\alpha_{k}+1$, whereas for $j=k$ the coefficient of $\partial_{k}F$ in $\mathbf{X}^{\boldsymbol{\alpha}}$ has weight $d-(\alpha_{k}+1)=(d-1)-\alpha_{k}$. We thus apply Corollary (1.12) to the homogeneity property (1.11), with $r_{j}=1$ for $j\neq k$ and $r_{k}=0$, which gives that $\mathrm{Res}\bigl(\partial_{0}F,\ldots,\partial_{n}F\bigr)$ is homogeneous of degree
$n(d-1)^{n+1}+n(d-1)^{n}.$
Again this gives the wanted result by (2.2.1). ∎
In turn, one may reproduce the argument given in (1.12) to deduce further weight properties from the two latter results and the standard homogeneity property of the discriminant stated in Definition (2.2).
(2.7) Corollary.
Assume that the discriminant $\mathrm{Disc}_{d}\in A_{\mathbf{Z}}$ is homogeneous of degree $\delta$ for the grading on $A_{\mathbf{Z}}$ defined by $\mathrm{weight}(U_{{\boldsymbol{\alpha}}})=w_{{\boldsymbol{\alpha}}}$. Let $r\in\mathbf{Z}$. For the new grading on $A_{\mathbf{Z}}$ defined by $\mathrm{weight}(U_{{\boldsymbol{\alpha}}})=w_{{\boldsymbol{\alpha}}}+r$, the discriminant $\mathrm{Disc}_{d}$ is homogeneous of degree
$\delta+r(n+1)(d-1)^{n}.$
The reduced discriminant
We write the generic homogeneous degree $d$ polynomial as
$F=X_{0}^{d}F_{0}+\cdots+X_{0}^{d-s}F_{s}+\cdots+F_{d}$
where each $F_{k}$ is homogeneous of degree $k$ in the indeterminates $X_{1},\ldots,X_{n}$. Given an integer $s\in[\kern-1.72218pt[2,d-1]\kern-1.72218pt]$, we consider the truncation $H$ of $F$ at order dsd-s with respect to $X_{0}$, defined as in Paragraph (1.2), and set $G$ as in Paragraph (1.2); thus,
$H$ $=$ $X_{0}^{d-s}F_{s}+X_{0}^{d-s-1}F_{s+1}+\cdots+F_{d},$
$G$ $=X_{0}^{s}F_{0}+X_{0}^{s-1}F_{1}+\cdots+$ $F_{s}.$
The truncation $H$ is the generic degree $d$ polynomial with a multiplicity $s$ zero at the point $(1\!:\!0\!:\!\ldots\!:\!0)$. The wish for the reduced discriminant is that it is a polynomial in the coefficients of $H$, the vanishing of which characterizes the existence of an additional singular non-trivial zero of $H$.
The polynomial $H$ is homogeneous of degree $d$, and has order $s$ in $X_{1},\ldots,X_{n}$ (recall Definition (0.2)). Moreover, its partial derivatives with respect to the indeterminates $X_{1},\ldots,X_{n}$ all have degree d1d-1, and order s1s-1 in $X_{1},\ldots,X_{n}$. Therefore, the reduced resultant
$\operatorname{redRes}(\partial_{1}H,\ldots,\partial_{n}H,H)=\operatorname{redRes}_{d-1,\ldots,d-1,d}^{s-1,\ldots,s-1,s}(\partial_{1}H,\ldots,\partial_{n}H,H)$
is well defined.
(2.2) Proposition.
With the above notation, the discriminants $\mathrm{Disc}(F_{s})$ and $\mathrm{Disc}(F_{d})$ both divide the reduced resultant $\operatorname{redRes}(\partial_{1}H,\ldots,\partial_{n}H,H)$.
Proof.
Both discriminants $\mathrm{Disc}(F_{s})$ and $\mathrm{Disc}(F_{d})$ are irreducible as elements of $A_{\mathbf{Z}}{}$, and therefore coprime because they do not depend on the same coefficients of $F$. To prove the proposition, it thus suffices to prove that the vanishing of $\mathrm{Disc}(F_{s})$ or $\mathrm{Disc}(F_{d})$ implies the vanishing of the reduced resultant. This follows from (1.4), as we shall now explain.
The vanishing of $\mathrm{Disc}(F_{d})$ at $f$ (a specialization of $F$ with coefficients in an algebraically closed field $\mathbf{k}$) implies the existence of a common root at infinity ($x_{0}=0$) of the polynomials $h,\partial_{1}h,\ldots,\partial_{n}h$ (with $h$ the specialization of $H$ corresponding to $f$). Indeed, if $\mathrm{Disc}(F_{d})$ vanishes at $f$, then by Theorem (2.4) there exists a non-trivial zero $\vphantom{x}{}^{\flat}\kern-1.00006ptx\in\mathbf{k}^{n}$ of $f_{d},\partial_{1}f_{d},\ldots\partial_{n}f_{d}$, and then $(0,\vphantom{x}^{\flat}\kern-1.00006ptx)\in\mathbf{k}^{n+1}$ is a non-trivial zero of $h,\partial_{1}h,\ldots,\partial_{n}h$, hence the reduced resultant vanishes at $f$ by (a) of (1.4).
Similarly, the vanishing of $\mathrm{Disc}(F_{s})$ at $f$ implies the existence of a common non-trivial zero of $h,\partial_{1}h,\ldots,\partial_{n}h$ infinitely near the point $p_{0}=(1\!:\!0\!:\!\ldots\!:\!0)$: indeed, $f_{s},\partial_{1}f_{s},\ldots,\partial_{n}f_{s}$ define the respective tangent cones at $p_{0}$ of the hypersurfaces $V(h),V(\partial_{1}h),\ldots,V(\partial_{n}h)$, and the vanishing of $\mathrm{Disc}(F_{s})$ at $f$ implies the existence of a non-trivial zero $\vphantom{x}{}^{\flat}\kern-1.00006ptx\in\mathbf{k}^{n}$ of $f_{s},\partial_{1}f_{s},\ldots\partial_{n}f_{s}$. Thus, by (b) of (1.4), the reduced resultant vanishes at $f$ if $\mathrm{Disc}(F_{s})$ vanishes at $f$. ∎
Alternatively, this proposition can be proved by means of inertia forms, as follows. By Theorem (1.3), the reduced resultant $\operatorname{redRes}(\partial_{1}H,\ldots,\partial_{n}H,H)$ belongs to the ideal of inertia forms $(\partial_{1}H,\ldots,\partial_{n}H,H):(X_{1},\ldots,X_{n})^{\infty}$. Therefore, for all integer $i=1,\ldots,n$, there exists an integer $N_{i}$ such that
(2.2.1) $X_{i}^{N_{i}}\operatorname{redRes}(\partial_{1}H,\ldots,\partial_{n}H,H)\in(\partial_{1}H,\ldots,\partial_{n}H,H);$
specializing the variable $X_{0}$ to 0 in (2.2.1), we get that
$X_{i}^{N_{i}}\operatorname{redRes}(\partial_{1}H,\ldots,\partial_{n}H,H)\in(\partial_{1}F_{d},\ldots,\partial_{n}F_{d},F_{d}).$
It follows that $\operatorname{redRes}(\partial_{1}H,\ldots,\partial_{n}H,H)$ belongs to the ideal of inertia forms $(\partial_{1}F_{d},\ldots,\partial_{n}F_{d},F_{d}):(X_{1},\ldots,X_{n})^{\infty}$. By Proposition (2.3), this ideal is generated by the discriminant of the polynomial $F_{d}$, hence $\mathrm{Disc}(F_{d})$ divides $\operatorname{redRes}(\partial_{1}H,\ldots,\partial_{n}H,H)$. A similar argument, albeit slightly more technical, can be used to show that $\mathrm{Disc}(F_{s})$ divides $\operatorname{redRes}(\partial_{1}H,\ldots,\partial_{n}H,H)$: see [16, Lemme I.1.3].
In the notation of (2.5), $F_{d}=\bar{F}$ and $F_{s}=\bar{G}$. Observe that $F_{d}$ and $F_{s}$ are generic homogeneous polynomials of degrees $d$ and $k$ respectively, in the indeterminates $(X_{1},\ldots,X_{n})$. Proposition (2.2) then leads to the following definition.
(2.3) Definition.
The reduced discriminant of $F$ with respect to the truncation at order dsd-s for the indeterminate $X_{0}$, denoted by $\operatorname{redDisc}^{s}_{d}(H)$, or simply $\operatorname{redDisc}(H)$, is defined by the equality
(2.3.1) $\mathrm{Disc}(F_{d})\,\mathrm{Disc}(F_{s})\operatorname{redDisc}(H)=\operatorname{redRes}\bigl(\partial_{1}H,\ldots,\partial_{n}H,H\bigr)\in A_{\mathbf{Z}}.$
For a degree $d$ polynomial $h\in\mathbf{k}[\mathbf{X}]$ with a singularity of order $s$ at the origin, $\operatorname{redDisc}_{d}^{s}(h)$ is defined as the specialization of $\operatorname{redDisc}_{d}^{s}(H)$ with respect to the unique specialization $A_{\mathbf{Z}}\to\mathbf{k}$ mapping $H$ to $h$.
Definition (2.3) implies that $\operatorname{redDisc}(H)$ is a primitive polynomial in $A_{\mathbf{Z}}{}$ (i.e., the greatest common divisor of its coefficients equals $1$), because $\operatorname{redRes}(\partial_{1}H,\ldots,\partial_{n}H,H)$ is primitive by [5, Proposition 4.24].
The identity (2.3.1) should be compared to (2.5.1). Beware that the reduced discriminant is not merely the reduced resultant of all the partial derivatives, because of the factor $\mathrm{Disc}(F_{s})$ in (2.3.1). The factor $\mathrm{Disc}(F_{d})$ is an artefact of our definition by considering $\operatorname{redRes}\bigl(\partial_{1}H,\ldots,\partial_{n}H,H\bigr)$ and not plainly $\operatorname{redRes}\bigl(\partial_{0}H,\partial_{1}H,\ldots,\partial_{n}H\bigr)$, in order to define the reduced discriminant as a primitive polynomial with integer coefficients without dealing with possible constant factors, similar to $d^{a(n,d)}$ in (2.2.1). The factor $\mathrm{Disc}(F_{s})$ on the other hand is geometrically meaningful, as we will see in (2.5) below.
According to (2.2.1), the classical discriminant of a homogeneous polynomial can be computed from the resultant of all its partial derivatives, up to the extraneous integer factor $d^{a(n,d)}$. It turns out that a similar formula holds for the reduced discriminant.
(2.4) Proposition.
The following identity in $A_{\mathbf{Z}}{}$ holds:
(2.4.1) $d^{a(n,d)-(s-1)^{n}}\mathrm{Disc}(F_{s})\operatorname{redDisc}(H)=\operatorname{redRes}(\partial_{0}H,\ldots,\partial_{n}H).$
Proof.
To prove this formula, we consider the reduced resultant $\mathcal{R}:=\operatorname{redRes}(\partial_{1}H,\ldots,\partial_{n}H,dH)$. First, from the homogeneity property (see Theorem (1.3)), we have
$\mathcal{R}:=d^{(d-1)^{n}-(s-1)^{n}}\operatorname{redRes}(\partial_{1}H,\ldots,\partial_{n}H,H).$
Next, the Euler Identity gives $\mathcal{R}:=\operatorname{redRes}(\partial_{1}H,\ldots,\partial_{n}H,\sum_{i=0}^{n}X_{i}\partial_{i}H)$, and we claim that
(2.4.2) $\mathcal{R}$ $=\operatorname{redRes}(\partial_{1}H,\ldots,\partial_{n}H,X_{0}\partial_{0}H)$
(2.4.3) $=(-1)^{n(d-1)^{n}}\mathrm{Res}(\partial_{1}F_{d},\ldots,\partial_{n}F_{d})\operatorname{redRes}(\partial_{1}H,\ldots,\partial_{n}H,\partial_{0}H)$
(observe that $\partial_{0}H$ has degree d1d-1 and order $s$ in $X_{1},\ldots,X_{n}$, and $\left.H\right|_{X_{0}=0}=F_{d}$). We conclude the proof of (2.4.1) before justifying the two above equalities. The Zariski formula together with the property of resultants under permutation of the polynomials [15, §5.8] yields the equality:
$\operatorname{redRes}(\partial_{1}H,\ldots,\partial_{n}H,\partial_{0}H)=(-1)^{n(d-1)^{n+1}-ns(s-1)^{n}}\operatorname{redRes}(\partial_{0}H,\ldots,\partial_{n}H).$
In addition, $\mathrm{Res}(\partial_{1}F_{d},\ldots,\partial_{n}F_{d})=d^{a(n-1,d)}\mathrm{Disc}(F_{d})$, so putting everything together, including (2.3.1), we deduce that
$d^{(d-1)^{n}-(s-1)^{n}-a(n-1,d)}\mathrm{Disc}(F_{s})\operatorname{redDisc}(H)=(-1)^{N}\operatorname{redRes}(\partial_{0}H,\ldots,\partial_{n}H),$
where $N=n(d-1)^{n}+n(d-1)^{n+1}-n(s-1)^{n}s=nd(d-1)^{n}-n(s-1)^{n}s$. Now, $N$ is an even integer, and $(d-1)^{n}-a(n-1,d)=a(n,d)$, hence the claimed formula.
To justify (2.4.2), we prove the following invariance property of the reduced resultant under some elementary transformations. Let $F_{1},\ldots,F_{n+1}$ be generic polynomials of degree $d-1,\ldots,d-1,d$ respectively, as well as generic linear forms $L_{1},\ldots,L_{n}$ in the variables $X_{1},\ldots,X_{n}$. Then, from the definition of the reduced resultant as a generator of the ideal of inertia forms in the generic setting, we deduce that
$\operatorname{redRes}_{d-1,\ldots,d-1,d}^{s-1,\ldots,s-1,s}(H_{1},\ldots,H_{n+1})\text{ divides }\operatorname{redRes}_{d-1,\ldots,d-1,d}^{s-1,\ldots,s-1,s}(H_{1},\ldots,H_{n+1}+\sum_{i=1}^{n}L_{i}H_{i}).$
We notice that this divisibility property remains valid under any specialization, in particular under the specialization sending $F_{n+1}$ to $F_{n+1}-\sum_{i=1}^{n}L_{i}F_{i}$, and leaving $F_{1},\ldots,F_{n}$ invariant. This implies that the above divisibility property also holds in the other direction, hence these two reduced resultant are equal, up to sign. Applying Zariski formula and the invariance of the resultant under elementary transformations [15, §5.9], we deduce that this sign equals 1.
The equality (2.4.3), on the other hand, is a consequence of the following multiplicativity property of the reduced resultant:
(2.4.4) $\operatorname{redRes}_{d_{1},\ldots,d_{n+1}+1}^{s_{1},\ldots,s_{n+1}}(H_{1},\ldots,H_{n},X_{0}H_{n+1})\\ =(-1)^{d_{1}\cdots d_{n}}\mathrm{Res}(\bar{F_{1}},\ldots,\bar{F_{n}})\cdot\operatorname{redRes}_{d_{1},\ldots,d_{n+1}}^{s_{1},\ldots,s_{n+1}}(H_{1},\ldots,H_{n},H_{n+1}),$
which follows from the Zariski formula. Indeed, let $F_{1},\ldots,F_{n+1}$ be generic polynomials as above, and denote by $H_{i}$ and $G_{i}$ their decompositions corresponding to the truncation at order $d_{i}-s_{i}$ for all $i$. On the one hand, we have (observe that $F_{n+1}$ and $X_{0}F_{n+1}$ have the same “GG”, namely $G_{n+1}$):
$\mathrm{Res}(F_{1},\ldots,F_{n},X_{0}F_{n+1})=\mathrm{Res}(G_{1},\ldots,G_{n},G_{n+1})\operatorname{redRes}(H_{1},\ldots,H_{n},X_{0}H_{n+1})\\ +\text{ terms of higher weight.}$
On the other hand, by properties of the classical resultant,
$\mathrm{Res}(F_{1},\ldots,F_{n},X_{0}F_{n+1})=(-1)^{d_{1}\ldots,d_{n}}\mathrm{Res}(\bar{F_{1}},\ldots,\bar{F_{n}})\mathrm{Res}(F_{1},\ldots,F_{n},F_{n+1}),$
hence, by the Zariski formula,
$\mathrm{Res}(F_{1},\ldots,F_{n},X_{0}F_{n+1})=\\ (-1)^{d_{1}\ldots,d_{n}}\mathrm{Res}(\bar{F_{1}},\ldots,\bar{F_{n}})\mathrm{Res}(G_{1},\ldots,G_{n+1})\operatorname{redRes}(H_{1},\ldots,H_{n+1})+\text{ terms of higher weight}.$
The comparison of the two above expressions of $\mathrm{Res}(F_{1},\ldots,F_{n},X_{0}F_{n+1})$ yields the expected equality. ∎
(2.5)Vanishing of the reduced discriminant
By definition of the truncation $H$, the generic hypersurface $\mathrm{V}(H)\subseteq\mathbf{P}^{n}$ has an ordinary $s$-fold point at the origin $(1\!:\!0\!:\!\ldots\!:\!0)$. By Paragraph (1.4), the vanishing of $\operatorname{redRes}\bigl(\partial_{1}h,\ldots,\partial_{n}h,h\bigr)$ for some specialization $h$ of $H$ in an algebraically closed field $\mathbf{k}$ corresponds to one of the two following properties:
  1. (a)
    the existence of a common zero for the polynomials $\partial_{1}h,\ldots,\partial_{n}h,h$ which is not the origin;
  2. (b)
    the existence of a common zero for the polynomials $\partial_{1}f_{s},\ldots,\partial_{n}f_{s},f_{s}$, equivalently a common root for the polynomials $\partial_{1}f_{s},\ldots,\partial_{n}f_{s}$ by Euler Identity, assuming for simplicity that the degree $d$ is non-zero in $\mathbf{k}$.
Property (b) is equivalent to the vanishing of $\mathrm{Disc}(f_{s})$, i.e., to the tangent cone of the hypersurface $\mathrm{V}(h)$ at the origin being a cone over a singular degree $s$ hypersurface (with the convention that a hypersurface of degree $s^{\prime}>s$ is a singular degree $s$ hypersurface); in other words, in the blow-up of $\mathrm{V}(h)$ at the origin, the exceptional divisor is singular. This is equivalent to the singularity of $V(h)$ at the origin being worse than an ordinary $s$-fold point, but does not imply that the blow-up of $V(h)$ is singular, see Example (2.6) below.
Property (a) is implied by the vanishing of $\mathrm{Disc}(f_{d})$, as we have seen in the proof of Proposition (2.2). The latter condition is equivalent to the hypersurface $\mathrm{V}(h)$ being tangent to the hyperplane at infinity $\mathrm{V}(X_{0})$.
In Definition (2.3) we discard the two factors $\mathrm{Disc}(F_{s})$ and $\mathrm{Disc}(F_{d})$; the upshot is that the vanishing of the reduced discriminant $\operatorname{redDisc}(H)$ defines an (irreducible444the proof of the irreducibility of the reduced discriminant will appear in [4], but this property also follows from the theory of toric discriminants (see [12, Chapter 9]), of which the reduced discriminants are particular instances. ) divisor which has as a dense subset the locus of those hypersurfaces $\mathrm{V}(h)$ that have a singularity off the origin. This implies, since smoothness is an open property, that if $\operatorname{redDisc}(H)$ vanishes at $h$, then the proper transform555here we take the following slightly abusive definition for the proper transform: let $\varepsilon:\tilde{\mathbf{P}}^{n}\to\mathbf{P}^{n}$ be the blow-up at the origin, with exceptional divisor $E$; the proper transform of $V(h)$ is the divisor $\varepsilon^{*}V(h)-sE$. If $V(h)$ has an ordinary $(s+1)$-fold at the origin, its genuine proper transform is $\hat{V}(h)=\varepsilon^{*}V(h)-(s+1)E$, which is smooth along $E$, whereas its proper transform in our sense is $\hat{V}(h)+E$, which is singular along $E$. of $V(h)$ in the blow-up of the origin is singular, in other words $V(h)$ has an additional singularity, possibly infinitely near, besides its $s$-fold point at the origin. In fact the converse holds as well, as we shall see in [4]. Thus, $\operatorname{redDisc}(H)$ vanishes at $h$ if and only $V(h)$ has an additional singularity, possibly infinitely near.
Note that we make a difference between having a singularity worse than an ordinary $s$-fold point and having an additional singularity infinitely near the ordinary $s$-fold point, see the following example and (2.8) below for more details.
(2.6) Example.
Set $n=2$ and $s=2$. In this case, $\mathrm{Disc}(F_{2})$ vanishes at $h$ if and only if $f_{2}$ is a square, equivalently, $h$ may be written as
$h=X_{0}^{d-2}\,X^{2}+X_{0}^{d-3}\,\bigl(a_{3}Y^{3}+a_{2}XY^{2}+a_{1}X^{2}Y+a_{0}X^{3}\bigr)+\,\cdots$
after a suitable change of coordinates $(X,Y)=\phi(X_{1},X_{2})$. Let $\varepsilon:\tilde{\mathbf{P}}^{2}\to\mathbf{P}^{2}$ be the blow-up at the origin. The relevant place to study the proper transform of $V(h)$ over the origin $(1\!:\!0\!:\!0)\in\mathbf{P}^{2}$ is the affine plane with coordinates $(U,Y)$ in $\tilde{\mathbf{P}}^{2}$ which dominates the affine plane $(X_{0}=1)$ in $\mathbf{P}^{2}$, and such that $\varepsilon^{*}(X,Y)=(UY,Y)$. In this affine chart, the pull-back $\varepsilon^{*}h$ is given by
$U^{2}Y^{2}+\bigl(a_{3}Y^{3}+a_{2}UY^{3}+a_{1}U^{2}Y^{3}+a_{0}U^{3}Y^{3}\bigr)+\,\cdots$
and the proper transform of $V(h)$ is defined by the equation obtained by factoring out $Y^{2}$ from $\varepsilon^{*}h$, i.e.,
$U^{2}+\bigl(a_{3}Y+a_{2}UY+a_{1}U^{2}Y+a_{0}U^{3}Y\bigr)+\,\cdots$
This shows that the proper transform of $V(h)$ is singular at some point lying on the exceptional divisor of $\varepsilon$ (defined by $Y=0$ in this affine chart) if and only if $a_{3}=0$.
The generic $h$ such that $\mathrm{Disc}(F_{2})$ vanishes defines a plane curve with an ordinary cusp at the origin (local equation $x^{2}+y^{3}=0)$; the proper transform of this curve in the blow-up at the origin is smooth (and tangent to the exceptional divisor), hence $\operatorname{redDisc}(H)$ does not vanish at $h$ even though $V(h)$ has a singularity worse than an ordinary double point at the origin. On the other hand, the generic $h$ such that $V(h)$ has an additional singularity infinitely near the origin (which implies that $\operatorname{redDisc}(H)$ vanishes) defines a plane curve with a tacnode at the origin (local equation $x^{2}+y^{4}=0$, two smooth local branches tangent to each other); the proper transform of such a curve has an ordinary double point located on the exceptional divisor.
(2.7) Remark.
In the reduced resultant situation, the condition that the $n+1$ hypersurfaces $\mathrm{V}(h_{1}),\ldots,\mathrm{V}(h_{n+1})$ have an infinitely near additional intersection point is not divisorial, i.e., it does not correspond to an irreducible element factoring out of $\operatorname{redRes}(H_{1},\ldots,H_{n+1})$. The reason is that this condition has codimension larger than one: indeed it amounts to the $n+1$ leading terms $f_{1,s_{1}},\ldots,f_{n+1,s_{n+1}}$ having a non-trivial common zero, in other words the $n+1$ corresponding hypersurfaces in $\mathbf{P}^{n-1}$ have non-trivial intersection: these are $n+1$ hypersurfaces in $\mathbf{P}^{n-1}$, hence the condition has codimension larger than one.
In contrast, in the reduced discriminant situation the condition that the hypersurface $\mathrm{V}(h)$ has a singularity worse than an ordinary $s$-fold point is divisorial, as it corresponds to the vanishing of $\mathrm{Disc}(F_{s})$ at the leading term $f_{s}$ of $h$. The vanishing of $\operatorname{redDisc}(H)$ at $h$ is not implied by the latter condition, but by the stronger condition that $V(h)$ has an additional singularity infinitely near its $s$-fold point.
We may use the Milnor number (see [I, Section **] for the definition) to give a precise meaning to the condition of having a singlarity worse than an ordinary $s$-fold point. Assume that $h$ defines a hypersurface with an isolated singularity at the origin, so that its Milnor number at the origin $\mu_{0}(h)$ is well-defined. If the singularity of $h$ at the origin is an ordinary $s$-fold point, then $\mu_{0}(h)=(s-1)^{n}$, and condition (b) in Paragraph (2.5) above is equivalent to the condition that $\mu_{0}(h)>(s-1)^{n}$ by [11, Theorem 1].
(2.9) Lemma.
In the standard grading of $A_{\mathbf{Z}}$, given by $\mathrm{weight}(U_{\boldsymbol{\alpha}})=1$ for all $\boldsymbol{\alpha}$}, the reduced discriminant $\operatorname{redDisc}_{d}^{s}$ is homogeneous of degree
(2.9.1) $(n+1)\bigl[(d-1)^{n}-(s-1)^{n}\bigr]-2n(s-1)^{n-1}.$
Proof.
This follows from the definition by a direct computation, since the degrees of the other quantities in Display (2.3.1) are known. It is arguably safer, however, to carry this out explicitly.
The two plain discriminants $\mathrm{Disc}(F_{d})$ and $\mathrm{Disc}(F_{s})$ are homogeneous of respective degrees $n(d-1)^{n-1}$ and $n(s-1)^{n-1}$, see Definition (2.2). Moreover, it follows from Theorem (1.3) that
$\operatorname{redRes}_{d-1,\ldots,d-1,d}^{s-1,\ldots,s-1,s}\,(\partial_{1}H,\ldots,\partial_{n}H,H)$
is homogeneous of degree
$n\bigl(d(d-1)^{n-1}-s(s-1)^{n-1}\bigr)+\bigl((d-1)^{n}-(s-1)^{n}\bigr).$
It thus follows from the equality in Display (2.3.1) that $\operatorname{redDisc}(H)$ is homogeneous, of degree
$\bigl[nd(d-1)^{n-1}+(d-1)^{n}-n(d-1)^{n-1}\bigr]-\bigl[ns(s-1)^{n-1}+(s-1)^{n}+n(s-1)^{n-1}\bigr]\\ =(n+1)(d-1)^{n}-(s-1)^{n-1}\bigl((n+1)s+n-1\bigr),$
and the result follows. ∎
We now turn to the generalized weight properties of the reduced discriminant.
(2.10) Example.
As an illustrative example, we consider the case $n=1$ and set
$F=\sum\nolimits_{i=0}^{d}U_{i}X_{0}^{d-i}X_{1}^{i}=X_{0}^{d}F_{0}+X_{0}^{d-1}F_{1}+\cdots+X_{0}^{d-s}F_{s}+\cdots F_{d},$
with $F_{i}=U_{i}X_{1}^{i}$ for all $i=0,\ldots,d$. We consider
$H=\sum\nolimits_{i=s}^{d}U_{i}X_{0}^{d-i}X_{1}^{i}=\sum\nolimits_{i=s}^{d}X_{0}^{d-i}F_{i}=X_{1}^{s}\bigl(U_{s}X_{0}^{d-s}+U_{s+1}X_{0}^{d-s-1}X_{1}+\cdots+U_{d}X_{1}^{d-s}\bigr),$
the truncation of $F$ at order ds>0d-s>0 with respect to $X_{0}$. Then, setting $H=X_{1}^{s}\cdot\vphantom{H}^{\flat}\kern-1.00006ptH$ it is easy to check that $\operatorname{redDisc}_{d}^{s}(H)=\pm\mathrm{Disc}(\vphantom{H}^{\flat}\kern-1.00006ptH)$; it is therefore an irreducible polynomial of degree 2(ds1)2(d-s-1) in the coefficients of $H$ (compare with Formula (2.9.1) for the degree).
Moreover, using the weight properties in Proposition (2.6), we can deduce weight properties of $\operatorname{redDisc}(H)$. Suppose that $A_{\mathbf{Z}}$ is graded with the rule $\mathrm{weight}(U_{i})=\max(0,i-s)$; then the reduced discriminant $\operatorname{redDisc}_{d}^{s}(H)$ is homogeneous of degree
(2.10.1) (ds)(ds1)(d-s)(d-s-1)
by Proposition (2.6), as $\vphantom{H}{}^{\flat}\kern-1.00006ptH$ has degree dsd-s. Similarly, if $A_{\mathbf{Z}}$ is graded with the rule $\mathrm{weight}(U_{i})=d-i$ then the same conclusion holds (note that di=(ds)(is)d-i=(d-s)-(i-s)).
We can generalize this following Corollary (1.12). Let $r$ be an integer and consider the grading of $A_{\mathbf{Z}}$ defined by the rule $\mathrm{weight}(U_{i})=i-s+r$ if $i\geqslant s$, and $\mathrm{weight}(U_{i})=0$ otherwise, then $\operatorname{redDisc}_{d}^{s}(H)$ is homogeneous of degree (ds+2r)(ds1)(d-s+2r)(d-s-1). In particular, if $r=s$, i.e., $\mathrm{weight}(U_{i})=i$ if $i\geqslant s$ and $\mathrm{weight}(U_{i})=0$ otherwise, we get that $\operatorname{redDisc}_{d}^{s}(H)$ is homogeneous of degree
(2.10.2) (d+s)(ds1)=d(d1)s(s+1).(d+s)(d-s-1)=d(d-1)-s(s+1).
In fact, since $\operatorname{redDisc}_{d}^{s}(H)$ a polynomial in $U_{s},\ldots,U_{d}$ only, the weights of $U_{0},\ldots,U_{s-1}$ don’t matter. Similarly, in the grading of $A_{\mathbf{Z}}$ is defined by the rule $\mathrm{weight}(U_{i})=d-i+r$ if $i\geqslant s$ and $\mathrm{weight}(U_{i})=0$ otherwise, $\operatorname{redDisc}_{d}^{s}(H)$ is homogeneous of degree (ds+2r)(ds1)(d-s+2r)(d-s-1).
The following result is similar to Theorem (1.3). It is the key to the generalized Salmon formula for the discriminant.
(2.11) Theorem ([5, Theorem 4.25]).
Suppose that the ring $A_{\mathbf{Z}}{}$ is graded by means of the Zariski grading (1.1), i.e., $\mathrm{weight}(U_{\boldsymbol{\alpha}})=\max(\alpha_{0}-d+s,0)$. Then $\mathrm{Disc}(F)$ has order $s(s-1)^{n}$, and its homogeneous part $D_{0}$ in this degree satisfies the following equality in $A_{\mathbf{Z}}$:
(2.11.1) $D_{0}\,\mathrm{Disc}(F_{d})=\mathrm{Disc}(G)\,\mathrm{Disc}(F_{s})\,\operatorname{redRes}\bigl(\partial_{1}H,\ldots,\partial_{n}H,H\bigr).$
Let us point out that the three elements $\mathrm{Disc}(F_{d})$, $\mathrm{Disc}(F_{s})$, and $\operatorname{redRes}(\partial_{1}H,\ldots,\partial_{n}H,H)$ have degree 0 with respect to the Zariski grading, while $\mathrm{Disc}(G)$ is homogeneous of degree $s(s-1)^{n}$ by (2.6.1); note that $F_{d}$ and $F_{s}$ are generic homogeneous polynomials of respective degrees $d$ and $s$ in the indeterminates $(X_{1},\ldots,X_{n})$, while $G$ is generic homogeneous of degree $s$ in $(X_{0},\ldots,X_{n})$.
(2.12) Remark.
The only trace towards the reduced discriminant in Salmon’s work that we have found is in [19, §117], where he proves, in the above language, that for $s=2$, $\mathrm{Disc}(F)$ has order at least $2$.
He obtains this as a direct consequence of the fact that if $f$ is a polynomial which is singular at a point $x$, then the tangent space of the discriminant hypersurface $\mathrm{V}(\mathrm{Disc})$ at $[f]$ contains the hyperplane of polynomials vanishing at $x$. Thus, for the polynomial
$f=\sum_{a>1}F_{a}X_{0}^{d-a}+T\cdot\bigl(U_{1}^{\prime}X_{1}+\cdots+U_{n}^{\prime}X_{n}\bigr)\,X_{0}^{d-1},$
one finds that $\mathrm{Disc}(f)$ is divisible by $T^{2}$ (here, $F$ is as in (2.1) and (2.1), and $T,U_{1}^{\prime},\ldots,U_{n}^{\prime}$ are new indeterminates). It follows that $\mathrm{Disc}(F)$ sits in the ideal
$\bigl(U_{d,0,\ldots,0}\bigr)+\bigl(U_{d-1,1,\ldots,0},\,\ldots\,,\,U_{d-1,0,\ldots,1}\bigr)^{2}\subseteq A_{\mathbf{Z}},$
which means that it has order at least $2$ with respect to the Zariski grading for $s=2$.
Proof of Theorem (2.11).
The idea is to specialize Theorem (1.3), part (a), and in particular the formula
$\bigl[\mathrm{Res}(F_{1},\ldots,F_{n+1})\bigr]_{s_{1}s_{2}\cdots s_{n+1}}=\mathrm{Res}(G_{1},\ldots,G_{n+1})\cdot\operatorname{redRes}(H_{1},\ldots,H_{n+1}),$
to the discriminant situation, i.e., $F_{1},\ldots,F_{n+1}$ specialize to $\partial_{1}F,\ldots,\partial_{n}F,F$ respectively, and the $G_{j}$’s and $H_{j}$’s specialize according to the truncations corresponding to $s_{1},\ldots,s_{n+1}=s-1,\ldots,s-1,s$, as indicated in Paragraph (2.1). Thus, we consider the following specializations:
$\mathrm{Res}(F_{1},\ldots,F_{n+1})$ $\leadsto\mathrm{Res}(\partial_{1}F,\ldots,\partial_{n}F,F)$
$\mathrm{Res}(G_{1},\ldots,G_{n+1})$ $\leadsto\mathrm{Res}(\partial_{1}G,\ldots,\partial_{n}G,G)$
$\operatorname{redRes}(H_{1},\ldots,H_{n+1})$ $\leadsto\operatorname{redRes}(\partial_{1}H,\ldots,\partial_{n}H,H)$
On the other hand, we have
($2.1$) $\mathrm{Res}(\partial_{1}F,\ldots,\partial_{n}F,F)$ $=\mathrm{Disc}(F)\,\mathrm{Disc}(\bar{F})$
(2.12.2) $\mathrm{Res}(\partial_{1}G,\ldots,\partial_{n}G,G)$ $=\mathrm{Disc}(G)\,\mathrm{Disc}(\bar{G})$
by Proposition (2.5). By definition, $\bar{F}=F_{d}$ and $\bar{G}=G_{s}=F_{s}$, and the result follows. We refer to [5] for more details. ∎
The following corollary is the general form of Salmon’s formula. As we will see in the next section, Salmon’s formula is a particular form of the corollary in the case $s=2$.
(2.13) Corollary.
In the Zariski grading of $A_{\mathbf{Z}}{}$ as in Theorem (2.11), the discriminant $\mathrm{Disc}(F)$ has order $s(s-1)^{n}$, and can be written as
(2.13.1) $\mathrm{Disc}(F)=\mathrm{Disc}(G)\,\mathrm{Disc}(F_{s})^{2}\,\operatorname{redDisc}(H)+\bigl(\mathrm{terms\ of\ Zariski\ weight}>s(s-1)^{n}\bigr).$
Proof.
The only novelty with respect to the previous theorem is the expression for $D_{0}$ (in the notation of Theorem (2.11)) in terms of the reduced discriminant. By Definition (2.3), one has
$\operatorname{redRes}\bigl(\partial_{1}H,\ldots,\partial_{n}H,H\bigr)=\mathrm{Disc}(F_{d})\,\mathrm{Disc}(F_{s})\,\operatorname{redDisc}(H).$
Thus, Equation (2.11.1) writes
$D_{0}\,\mathrm{Disc}(F_{d})=\mathrm{Disc}(G)\,\mathrm{Disc}(F_{s})\,\mathrm{Disc}(F_{d})\,\mathrm{Disc}(F_{s})\,\operatorname{redDisc}(H)$
which gives
$D_{0}=\mathrm{Disc}(G)\,\mathrm{Disc}(F_{s})^{2}\operatorname{redDisc}(H)$
as we wanted. ∎
We point out that, in the above, the factor $\mathrm{Disc}(F_{s})$ comes once from Formula (2.11.1), where it came from Formula (2.12.2) as $\mathrm{Disc}(\bar{G})$, and once from the definition of $\operatorname{redDisc}(H)$, which explains the square in Formula (2.13.1). In the discussion following Definition (2.3) above, we explained that, in that definition, $\mathrm{Disc}(F_{s})$ is a geometrically meaningful term, as $F_{s}$ is the equation of the tangent cone of $\mathrm{V}(H)$ at the origin, whereas $\mathrm{Disc}(F_{d})$ is a technical artefact. It is thus fortunate that the latter term does not appear in Formula (2.13.1). Note that, indeed, the factor $\mathrm{Disc}(F_{d})$ in the definition of $\operatorname{redDisc}(H)$ cancels out with the same factor coming from Formula (2.12.1) as $\mathrm{Disc}(\bar{F})$.
Corollary (2.13) provides an interesting connection between ordinary and reduced discriminants. As a first illustration of its interest, we give the following generalized weight properties of the reduced discriminant, which generalize the computations of Example (2.10) to arbitrary $n$.
(2.14) Proposition.
In the grading of $A_{\mathbf{Z}}$ defined by $\mathrm{weight}(U_{\boldsymbol{\alpha}})=\alpha_{0}$, the reduced discriminant $\operatorname{redDisc}_{d}^{s}(F)$ is homogeneous of degree
$d(d-1)^{n}-(s-1)^{n-1}\bigl[d\,\bigl((n+1)s+n-1\bigr)-ns(s+1)\bigr].$
In the grading of $A_{\mathbf{Z}}$ defined by $\mathrm{weight}(U_{\boldsymbol{\alpha}})=d-\alpha_{0}$, the reduced discriminant $\operatorname{redDisc}_{d}^{s}(F)$ is homogeneous of degree
$n\bigl[d(d-1)^{n}-s(s+1)(s-1)^{n-1}\bigr].$
Proof.
In the two gradings of (2.14.1) and (2.14.2), the plain discriminant $\mathrm{Disc}(F)$ is homogeneous. Thus all its summands are homogeneous, and in particular its piece of lowest degree in the Zariski grading is homogeneous. We shall see that $\mathrm{Disc}(G)$ and $\mathrm{Disc}(F_{s})$ are homogeneous as well. Therefore, it follows from Formula (2.13.1) that $\operatorname{redDisc}(H)$ is homogeneous. Then it is only a matter of computing its degree.
In the grading of (2.14.1), $\mathrm{Disc}_{d}(F)$ has weight $d(d-1)^{n}$ by Proposition (2.6). One has $G=X_{0}^{s}F_{0}+X_{0}^{s-1}F_{1}+\cdots+F_{s}$ and the coefficients of $F_{i}$ have weight di=(si)+(ds)d-i=(s-i)+(d-s), hence $\mathrm{Disc}_{d}(G)$ has weight
$s(s-1)^{n}+(n+1)(d-s)(s-1)^{n}$
by Proposition (2.6) and Corollary (2.7). The coefficients of $F_{s}$ all have weight dsd-s, hence $\mathrm{Disc}_{s}(F_{s})$ has weight
$n(d-s)(s-1)^{n-1}$
by the homogeneity property of the discriminant in Definition (2.2) and Corollary (2.7) (note that $\mathrm{Disc}_{s}(F_{s})$ is a discriminant for polynomials in $n$ indeterminates, i.e., one less than $\mathrm{Disc}_{d}(F)$ and $\mathrm{Disc}_{d}(G)$). One thus finds, by the above Corollary (2.13), that $\operatorname{redDisc}_{d}^{s}(F)$ is homogeneous of degree
$d(d-1)^{n}-s(s-1)^{n}-(n+1)(d-s)(s-1)^{n}-2n(d-s)(s-1)^{n-1}\\ \begin{aligned} &=d(d-1)^{n}-(s-1)^{n-1}\bigl[s(s-1)+(n+1)(d-s)(s-1)+2n(d-s)\bigr]\\ &=d(d-1)^{n}-(s-1)^{n-1}\bigl[d\,\bigl((n+1)s+(n-1)\bigr)-ns(s+1)\bigr].\end{aligned}$
In the grading of (2.14.2), $\mathrm{Disc}(F)$ is homogeneous of degree $nd(d-1)^{n}$ by Proposition (2.6), and $\mathrm{Disc}(G)$ is homogeneous of degree $ns(s-1)^{n}$ by the same result. In addition, all the coefficients of $F_{s}$ have weight $s$ in this grading hence, by the homogeneity property given in Definition (2.2), $\mathrm{Disc}(F_{s})$ is homogeneous of degree $ns(s-1)^{n-1}$ (note that $F$ is a homogeneous polynomial in $n$ variables only). Then the conclusion follows by a direct computation. ∎
For $n=1$, the degree in (2.14.1) is
$d(d-1)-\bigl[2sd-s(s+1)\bigr]=d^{2}-(2s+1)d+s(s+1),$
which agrees with (2.10.1), and the degree in (2.14.2) agrees with (2.10.2).
In the above paragraphs we have introduced the reduced discriminant and provided some first properties that are sufficient for our purposes. A more detailed and complete study of this new eliminant polynomial, including for instance its irreducibility and the geometric meaning of its vanishing, in particular its connection to Milnor number, will appear in [4].
Application to the Salmon Formula
We shall now see that Salmon’s formula (a) is a particular case of the decomposition formula given in Corollary (2.13). We will then be able to generalize it to the case of a hypersurface in arbitrary dimension.
(2.1)Original Salmon Formula
Let $F(X,Y,Z)$ be the generic homogeneous polynomial of degree $d\geqslant 3$, and set
$F=X^{d}F_{0}+X^{d-1}F_{1}+X^{d-2}F_{2}+\cdots+XF_{d-1}+F_{d},$
$F_{0}=U,\qquad F_{1}(Y,Z)=SY+TZ,\qquad F_{2}(Y,Z)=\frac{1}{2}(A{Y}^{2}+2BYZ+C{Z}^{2});$
(we introduce a rational number in the definition of $F_{2}$ to follow Salmon’s notation; this is natural in terms of the Taylor–Newton Formula of [II, Paragraph **]).
We consider the truncation of $F$ at order d2d-2 with respect to $X$, and the corresponding Zariski grading. So, $U$ has degree 2, $S$ and $T$ have degree 1, and the coefficients of the $F_{k}$’s with $k\geqslant 2$ have degree 0 (this includes the coefficients $A,B,C$). We let
$H=X^{d-2}F_{2}+\cdots+XF_{d-1}+F_{d}\quad\text{and}\quad G=X^{2}F_{0}+XF_{1}+F_{2}.$
Then, Corollary (2.13) tells us that
(2.1.1) $\mathrm{Disc}(F)=\mathrm{Disc}(G)\,\mathrm{Disc}(F_{2})^{2}\,\operatorname{redDisc}(H)+\bigl(\text{terms of Zariski weight}\geqslant 3\bigr).$
Discriminants of quadratic forms are readily computed, see, e.g., [8, no5 Exemple 6, p. 363], which is based on Identity (2.2.1); one has
$\mathrm{Disc}(F_{2})=\begin{vmatrix}A&B\\ B&C\end{vmatrix}=AC-B^{2},$
and
$\mathrm{Disc}(G)=\frac{1}{2}\cdot\begin{vmatrix}2U&S&T\\ S&A&B\\ T&B&C\end{vmatrix}=\frac{1}{2}\left(2ACU-A{T}^{2}-2{B}^{2}{U}+2BST-C{S}^{2}\right).$
Thus, Display (2.1.1) reads
$\mathrm{Disc}(F)=\frac{1}{2}\left(2ACU-A{T}^{2}-2{B}^{2}{U}+2BST-C{S}^{2}\right)\cdot\left(AC-B^{2}\right)^{2}\cdot\operatorname{redDisc}(H)\\ \ \textrm{mod}\ \bigl((S,T)^{3}+U(S,T)+(U^{2})\bigr).$
Then, the specialization $U=S=0$ yields
(2.1.2) $\mathrm{Disc}(F)=-\frac{1}{2}A\,{T}^{2}\cdot\bigl(AC-B^{2}\bigr)^{2}\cdot\operatorname{redDisc}(H)\quad\ \textrm{mod}\ (T^{3});$
this is Salmon’s Formula (a), with the normalization factor $\frac{1}{2}$, and with $\Phi=\operatorname{redDisc}(H)$.
(2.2)Salmon Formula in arbitrary dimension
The Salmon Formula for the discriminant of a plane curve can be generalized to the case of a hypersurface in a projective space of arbitrary dimension as follows. In a suitable system of homogeneous coordinates, any hypersurface $\mathrm{V}(F)\subseteq\mathbf{P}^{n}$ has an equation of the form
(2.2.1) $F(X_{0},X_{1},\ldots,X_{n})=TX_{0}^{d-1}X_{n}+\sum_{k=2}^{d}X_{0}^{d-k}F_{k}(X_{1},\ldots,X_{n}),$
where for all $k=2,\ldots,d$, the polynomial $F_{k}$ is homogeneous of degree $k$ in the indeterminates $X_{1},\ldots,X_{n}$. This normal form merely imposes that the hypersurface $\mathrm{V}(F)$ goes through the point $(1:0:\cdots:0)$ and that its tangent hyperplane at this point is given by $X_{n}=0$. Applying Corollary (2.13) as above, and setting $H=\sum_{k=2}^{d}X_{0}^{d-k}F_{k}$, we find that
$\mathrm{Disc}(F)=\mathrm{Disc}(TX_{0}X_{n}+F_{2})\,\mathrm{Disc}(F_{2})^{2}\,\operatorname{redDisc}(H)\quad\ \textrm{mod}\ T^{3}.$
Let $\bar{F}_{2}(X_{1},\ldots,X_{n-1})$ be the homogeneous polynomial of degree $2$ in the indeterminates $X_{1},\ldots,X_{n-1}$ defined as
$\bar{F}_{2}(X_{1},\ldots,X_{n-1})=F_{2}(X_{1},\ldots,X_{n-1},0)$
(beware the difference in notation with (2.5)). We have:
$2^{a(n,2)}\,\mathrm{Disc}(TX_{0}X_{n}+F_{2})$
$=\mathrm{Res}\bigl(\partial_{0}(TX_{0}X_{n}+F_{2}),\,\partial_{1}(TX_{0}X_{n}+F_{2}),\,\ldots,\,\partial_{n}(TX_{0}X_{n}+F_{2})\bigr)$
$=\mathrm{Res}\bigl(TX_{n},\,\partial_{1}F_{2},\,\ldots,\,\partial_{n-1}F_{2},\,TX_{0}+\partial_{n}F_{2}\bigr)$
(2.2.2) $=-\mathrm{Res}\bigl(TX_{0}+\partial_{n}F_{2},\,\partial_{1}F_{2},\,\ldots,\,\partial_{n-1}F_{2},\,TX_{n}\bigr)$
(2.2.3) $=-T\cdot\mathrm{Res}\bigl(TX_{0}+\left.\partial_{n}F_{2}\right|_{X_{n}=0},\,\left.\partial_{1}F_{2}\right|_{X_{n}=0},\,\ldots,\,\left.\partial_{n-1}F_{2}\right|_{X_{n}=0}\bigr)$
(2.2.4) $=-T^{2}\cdot\mathrm{Res}(\partial_{1}\bar{F}_{2},\,\ldots,\,\partial_{n-1}\bar{F}_{2})=-2^{a(n-2,2)}\,T^{2}\cdot\mathrm{Disc}(\bar{F}_{2}),$
where:
  • the exponents $a(n,2)$ and a(n2,2)a(n-2,2) are as defined in (2.2);
  • (2.2.2) is obtained by [15, §5.8, Effet d’une permutation des polynômes];
  • (2.2.3) is obtained by [15, (5.13.5)];
  • (2.2.4) is obtained by [15, §5.10, Formule de Laplace].
Alternatively, since the resultants above are resultants of homogeneous polynomials of degree $1$ they may be computed as determinants, see [15, §5.3, Cas des formes linéaires], to the effect that
(2.2.5) $\mathrm{Res}\bigl(TX_{n},\,\partial_{1}F_{2},\,\ldots,\,\partial_{n-1}F_{2},\,TX_{0}+\partial_{n}F_{2}\bigr)=\begin{vmatrix}0&0&\cdots&\cdots&0&T\\ 0&[\phantom{A}&&\partial_{1}F_{2}&&\phantom{A}]\\ \vdots&&&\vdots&\\ 0&[\phantom{A}&&\partial_{n-1}F_{2}&&\phantom{A}]\\ T&[\phantom{A}&&\partial_{n}F_{2}&&\phantom{A}]\end{vmatrix}$,
[…]
Either way, the upshot is that $\mathrm{Disc}(TX_{0}X_{n}+F_{2})=-T^{2}\,\mathrm{Disc}(\bar{F}_{2})$ (note that a(n,2)=a(n2,2)a(n,2)=a(n-2,2)), and eventually we obtain the following generalized Salmon formula:
(2.2.6) $\mathrm{Disc}(F)=-T^{2}\,\mathrm{Disc}(\bar{F}_{2})\,\mathrm{Disc}(F_{2})^{2}\operatorname{redDisc}(H)\quad\ \textrm{mod}\ T^{3}.$
In the situation of the original Salmon Formula above, one has $F_{2}(Y,Z)=\frac{1}{2}(AY^{2}+2BYZ+CZ^{2})$ and $\bar{F}_{2}(Y)=\frac{1}{2}AY^{2}$, hence $\mathrm{Disc}(F_{2})=AC-B^{2}$ and $\mathrm{Disc}(\bar{F}_{2})=\frac{1}{2}A$, and thus the two formulas (2.1.2) and (2.2.6) are coherent.
Computation of the node-couple degree by elimination
In [18, §605–607], Salmon sets up the following strategy to compute the number of $2$-nodal curves in a general net of hyperplane sections of a smooth (hyper)surface $S\subseteq\mathbf{P}^{3}$. For $p^{\prime}\in S$ and $p^{\prime\prime}\in\mathbf{T}_{p^{\prime}}S-\{p^{\prime}\}$, consider the pencil $\langle p^{\prime},p^{\prime\prime}\rangle^{\perp}\subseteq\check{\mathbf{P}}^{3}$ of (hyper)planes containing $p^{\prime}$ and $p^{\prime\prime}$. It cuts out $d^{\vee}=\deg(S^{\vee})$ points on the dual surface $S^{\vee}$, counted with multiplicities, which correspond to planes tangent to $S$. Among these, $\mathbf{T}_{p^{\prime}}S$ counts doubly if it is a plain tangent plane, and triply if it is plainly bitangent.
Indeed, the line $\langle p^{\prime},p^{\prime\prime}\rangle^{\perp}$ is contained in the plane $(p^{\prime})^{\perp}$, hence tangent to $S^{\vee}$ at the point $(\mathbf{T}_{p^{\prime}}S)^{\perp}\in S^{\vee}$. If the plane $\mathbf{T}_{p^{\prime}}S$ is plainly bitangent to $S$, then $(\mathbf{T}_{p^{\prime}}S)^{\perp}$ is a general point on the ordinary double curve of $S^{\vee}$, and the line $\langle p^{\prime},p^{\prime\prime}\rangle^{\perp}$ is tangent to one of the two transverse sheets of $S^{\vee}$ at $(\mathbf{T}_{p^{\prime}}S)^{\perp}$. The idea is then to determine first the conditions on $p^{\prime}$ for $\mathbf{T}_{p^{\prime}}S$ to count with multiplicity greater than $2$ in $\langle p^{\prime},p^{\prime\prime}\rangle^{\perp}$, and then to sort out the various corresponding geometric situations. A key element to carry this out is the famous formula (a); another one is the elimination procedure (3.2).
We work out Salmon’s procedure in subsection 3.2, and in subsection 3.3 we show how it carries over for hypersurfaces in a projective space of arbitrary dimension. From now on, we work over an algebraically closed field $\mathbf{k}$ of characteristic 0. We use freely the theory of polarity, on which there is a recap in the previous chapter [II, Appendix A].
An elimination trick
This section is dedicated to an elimination trick due to Salmon, see [18, §606]. It is the main technical device that he uses to conduct his study of pencils of planes orthogonal to a tangent line, which eventually gives the node-couple degree. We advise the reader to skip this section in first reading and move on to Section 3.2 where the general picture of Salmon’s approach is given, and to come back here only when needed.
Let $\mathbf{k}$ be an algebraically closed field of characteristic 0, and consider a smooth, degree $d$, surface $S\subseteq\mathbf{P}^{3}$ defined by a homogeneous polynomial $F\in\mathbf{k}[X,Y,Z,W]$; set
$\mathcal{T}_{S}=\bigl\{(p^{\prime},p^{\prime\prime}):p^{\prime}\in S\text{ and }p^{\prime\prime}\in\mathbf{T}_{p^{\prime}}S\bigr\}\,\subseteq\,\mathbf{P}^{3}\times\mathbf{P}^{3}.$
[…]
(3.2) Theorem.
Let $p^{\prime}$ be a point of $S$. The following assertions are equivalent:
  1. (i)
    there exists $p^{\prime\prime}\in\mathbf{T}_{p^{\prime}}S-\{p^{\prime}\}$ such that the tangent plane $\mathbf{T}_{p^{\prime}}S$ is a component of $\mathrm{V}(\Phi_{\hat{p}^{\prime},\hat{p}^{\prime\prime}})$;
  2. (ii)
    for all $p^{\prime\prime}\in\mathbf{T}_{p^{\prime}}S-\{p^{\prime}\}$, $\mathbf{T}_{p^{\prime}}S$ is a component of $\mathrm{V}(\Phi_{\hat{p}^{\prime},\hat{p}^{\prime\prime}})$.
Moreover, there exists a homogeneous polynomial $G\in\mathbf{k}[X,Y,Z,W]$ of degree $\lambda+(d-2)\mu$ such that (i) and (ii) are equivalent to:
  1. (iii)
    $p^{\prime}$ lies on the hypersurface $\mathrm{V}(G)\subseteq\mathbf{P}^{3}$.
We need some preparation for the proof. We begin with the following characterization of the non-emptiness of the intersection of two lines in $\mathbf{P}^{3}$, one given parametrically and the other defined by equations.
(3.3) Lemma.
Let $L^{\prime},L^{\prime\prime}\in\mathbf{k}[X,Y,Z,W]$ be two linear functionals, and $\hat{p}^{\prime},\hat{p}^{\prime\prime}\in\mathbf{k}^{4}$. The intersection of the two lines $\mathrm{V}(L^{\prime},L^{\prime\prime})$ and $\langle p^{\prime},p^{\prime\prime}\rangle$ is non-empty if and only if
$D_{L^{\prime},L^{\prime\prime}}(p^{\prime},p^{\prime\prime})\overset{\text{def.}}{=}\begin{vmatrix}L^{\prime}(\hat{p}^{\prime})&L^{\prime\prime}(\hat{p}^{\prime})\\ L^{\prime}(\hat{p}^{\prime\prime})&L^{\prime\prime}(\hat{p}^{\prime\prime})\end{vmatrix}$
vanishes.
Proof.
The line $\langle p^{\prime},p^{\prime\prime}\rangle$ is the image of the map $\mathbf{P}^{1}\rightarrow\mathbf{P}^{3}$ defined by the linear map $(u,v)\in\mathbf{k}^{2}\mapsto u\hat{p}^{\prime}+v\hat{p}^{\prime\prime}\in\mathbf{k}^{4}$. Therefore, the lines $\mathrm{V}(L^{\prime},L^{\prime\prime})$ and $\langle p^{\prime},p^{\prime\prime}\rangle$ intersect if and only if the two polynomials $L^{\prime}(u\hat{p}^{\prime}+v\hat{p}^{\prime\prime})$ and $L^{\prime\prime}(u\hat{p}^{\prime}+v\hat{p}^{\prime\prime})$ share a common root in $\mathbf{P}^{1}$. These polynomials are linear forms in $u,v$, and the polynomial $D_{L^{\prime},L^{\prime\prime}}(p^{\prime},p^{\prime\prime})$ is the determinant of the corresponding linear system. ∎
For all $L^{\prime},L^{\prime\prime}$, $D_{L^{\prime},L^{\prime\prime}}$ is a bihomogeneous polynomial in the two sets of variables $(p^{\prime},p^{\prime\prime})$, of bidegree $(1,1)$ and anti-symmetric; this implies that it is irreducible.
(3.4) Lemma.
For general linear forms $L^{\prime},L^{\prime\prime}$, the intersection in $\mathbf{P}^{3}\times\mathbf{P}^{3}$ of $\mathcal{T}_{S}$ and the hypersurface $\mathrm{V}(D_{L^{\prime},L^{\prime\prime}})$ is integral.
Proof.
A natural idea is to try and prove the lemma as an application of Bertini’s theorem. It is however delicate to proceed this way, as for instance the map defined on $\mathbf{P}^{3}\times\mathbf{P}^{3}$ by the linear system of anti-symmetric forms of bidegree $(1,1)$ is certainly not generically finite, and even less so the map defined by the system of all $D_{L^{\prime},L^{\prime\prime}}$. We will thus rather prove the result by hand.
Let us first note that as $L^{\prime},L^{\prime\prime}$ range through all linear forms, the bihomogeneous forms $D_{L^{\prime},L^{\prime\prime}}$ cover all rank $2$, bilinear, anti-symmetric forms. This is a direct calculation: for
$L^{\prime}=A^{\prime}X+B^{\prime}Y+C^{\prime}Z+D^{\prime}W\quad\text{and}\quad L^{\prime\prime}=A^{\prime\prime}X+B^{\prime\prime}Y+C^{\prime\prime}Z+D^{\prime\prime}W,$
$D_{L^{\prime},L^{\prime\prime}}$ is the bilinear form
$\left(\begin{pmatrix}X^{\prime}\\ Y^{\prime}\\ Z^{\prime}\\ W^{\prime}\end{pmatrix},\begin{pmatrix}X^{\prime\prime}\\ Y^{\prime\prime}\\ Z^{\prime\prime}\\ W^{\prime\prime}\end{pmatrix}\right)\longmapsto(X^{\prime},Y^{\prime},Z^{\prime},W^{\prime})\begin{pmatrix}0&U_{01}&U_{02}&U_{03}\\ -U_{01}&0&U_{12}&U_{13}\\ -U_{02}&-U_{12}&0&U_{23}\\ -U_{03}&-U_{13}&-U_{23}&0\end{pmatrix}\begin{pmatrix}X^{\prime\prime}\\ Y^{\prime\prime}\\ Z^{\prime\prime}\\ W^{\prime\prime}\end{pmatrix}$
where the coefficients $U_{ij}$ are the six $2\times 2$ minors of the matrix
$\begin{pmatrix}A^{\prime}&B^{\prime}&C^{\prime}&D^{\prime}\\ A^{\prime\prime}&B^{\prime\prime}&C^{\prime\prime}&D^{\prime\prime}\end{pmatrix}$.
It follows that the only relation between the coefficients $U_{ij}$ is the vanishing of the Pfaffian of the anti-symmetric matrix $(U_{ij})$, i.e.,
$U_{01}U_{23}-U_{02}U_{13}+U_{03}U_{12}=0$
(this is the Plücker equation of the Grassmannian $\mathrm{Gr}(2,4)$ in $\mathbf{P}^{5}$).
Now to the lemma. Projection on the first factor in $\mathbf{P}^{3}\times\mathbf{P}^{3}$ makes $\mathcal{T}_{S}$ a $\mathbf{P}^{2}$ bundle over $S$: for all $p^{\prime}\in S$, the fibre over $p^{\prime}$ is the zero locus of the linear form $\mathrm{D}_{p^{\prime}}F$ (the differential of $F$, a defining equation for $S$, at the point $p^{\prime}$; it is non-zero as we assume that $S$ is smooth). Similarly, for all $p^{\prime}\in\mathbf{P}^{3}$ the fibre $\mathrm{pr}_{1}^{-1}(p^{\prime})\cap\mathrm{V}(D_{L^{\prime},L^{\prime\prime}})$ is the zero locus of the linear form $D_{L^{\prime},L^{\prime\prime}}(p^{\prime},\,\_\,)$, which is either a hyperplane or the whole $\mathbf{P}^{3}$. We claim that for general $L^{\prime},L^{\prime\prime}$, the intersection $\mathcal{T}_{S}\cap\mathrm{V}(D_{L^{\prime},L^{\prime\prime}})$ is the closure of a $\mathbf{P}^{1}$-bundle over an open subset of $S$, which implies at once that it is irreducible. Moreover, this implies that the intersection $\mathcal{T}_{S}\cap\mathrm{V}(D_{L^{\prime},L^{\prime\prime}})$ is fibrewise over $S$ a transverse intersection of hyperplanes, and therefore it is reduced. Thus the claim proves the lemma. To prove the claim it suffices to show that, for general $L^{\prime},L^{\prime\prime}$, the set of points $p^{\prime}\in S$ such that the linear form $D_{L^{\prime},L^{\prime\prime}}(p^{\prime},\,\_\,)$ is a multiple of $\mathrm{D}_{p^{\prime}}F$ has dimension 0; in particular, the intersection $\mathcal{T}_{S}\cap\mathrm{V}(D_{L^{\prime},L^{\prime\prime}})$ cannot have any irreducible component generically a $\mathbf{P}^{2}$-bundle over a curve in $S$.
The claim follows from explicit computations. The two linear forms $\mathrm{D}_{p^{\prime}}(F)$ and $D_{L^{\prime},L^{\prime\prime}}(p^{\prime},\,\_\,)$ are, in coordinates,
$\bigl(\partial_{X}F(p^{\prime}),\,\partial_{Y}F(p^{\prime}),\,\partial_{Z}F(p^{\prime}),\,\partial_{W}F(p^{\prime})\bigr)\quad\text{and}\quad\begin{pmatrix}X^{\prime\prime}\\ Y^{\prime\prime}\\ Z^{\prime\prime}\\ W^{\prime\prime}\end{pmatrix}^{\kern-2.45836pt\mathsf{T}}\cdot\begin{pmatrix}0&U_{01}&U_{02}&U_{03}\\ -U_{01}&0&U_{12}&U_{13}\\ -U_{02}&-U_{12}&0&U_{23}\\ -U_{03}&-U_{13}&-U_{23}&0\end{pmatrix}$.
Both have the line $p^{\prime}\subseteq\mathbf{k}^{4}$ in their kernel, so we consider the corresponding linear forms on the quotient $\mathbf{k}^{4}\big/p^{\prime}$, given by line matrices with three entries; up to a linear change of coordinates we may assume that $p^{\prime}=(0:0:0:1)$, in which case we are simply forgetting the last entry in the two above line matrices. The claim then boils down to the fact that the three $2\times 2$ minors of the $2\times 3$ matrix obtained by concatenating the two line matrices corresponding to $\mathrm{D}_{p^{\prime}}(F)$ and $D_{L^{\prime},L^{\prime\prime}}(p^{\prime},\,\_\,)$ define a locus of codimension at least two on $S$, for general $L^{\prime},L^{\prime\prime}$. Since everything is explicit, one may observe that for a general choice of $L^{\prime},L^{\prime\prime}$, any two of these three minors have no common factor, which proves the claim. ∎
(3.5)Proof of Theorem (3.2)
The proof is based on the fact that $\mathbf{T}_{p^{\prime}}S$ is a component of $\mathrm{V}(\Phi_{p^{\prime},p^{\prime\prime}})$ if and only if, for all line $\Lambda\subseteq\mathbf{P}^{3}$ the intersection $L\cap\mathbf{T}_{p^{\prime}}S\cap\mathrm{V}(\Phi_{p^{\prime},p^{\prime\prime}})$ is non-empty; a moment of thought should convince the reader that this fact is indeed true. Then the idea is to consider a suitable resultant to express this fact, and to factor a power of $D_{L^{\prime},L^{\prime\prime}}(p^{\prime},p^{\prime\prime})$ out of this resultant.
Consider $T_{p^{\prime}}=\mathrm{D}^{p}F(p^{\prime})$ as a trihomogeneous polynomial of tridegree (d1,0,1)(d-1,0,1) in $(p^{\prime},p^{\prime\prime},p)$. Let $L^{\prime},L^{\prime\prime}\in\check{\mathbf{P}}^{3}$ be two generic linear forms in $p$, and consider the line $\Lambda=\mathrm{V}(L^{\prime},L^{\prime\prime})$ in $\mathbf{P}^{3}$. It intersects $\mathbf{T}_{p^{\prime}}S\cap\mathrm{V}(\Phi_{p^{\prime},p^{\prime\prime}})$ if and only if the resultant
$R=\mathrm{Res}(T_{p^{\prime}},\Phi_{p^{\prime},p^{\prime\prime}},L^{\prime},L^{\prime\prime}),$
with respect to the set of variables $p$, vanishes. It follows from the homogeneity properties of the resultant, see Theorem (1.7), that $R$ is a multi-homogeneous polynomial in the sets of variables $(L^{\prime},L^{\prime\prime},p^{\prime},p^{\prime\prime})$, of multi-degree $(\mu,\mu,\mu(d-1)+\lambda,\mu)$.
For a given $(p^{\prime},p^{\prime\prime})\in\mathcal{T}_{S}$, the polynomial $\Phi_{p^{\prime},p^{\prime\prime}}$ splits as the product of $\mu$ linear forms $\Phi^{i}_{p^{\prime},p^{\prime\prime}}$, $i=1,\ldots,\mu$, that all vanish along the line $\langle p^{\prime},p^{\prime\prime}\rangle$. Although this splitting may not exist globally over $\mathcal{T}_{S}$ (in other words the factors $\Phi^{i}_{p^{\prime},p^{\prime\prime}}$ may not be polynomials in $(p^{\prime},p^{\prime\prime})$), there exists a global splitting over a finite cover $\tilde{\mathcal{T}_{S}}$ of $\mathcal{T}_{S}$666let us briefly recap how such a finite cover may be constructed. The family of planes defined over $\mathcal{T}_{S}$ by $\Phi_{p^{\prime},p^{\prime\prime}}$ corresponds to a subscheme $\mathcal{P}$ of $\mathcal{T}_{S}\times\check{\mathbf{P}^{3}}$ with the first projection inducing a degree $\mu$ map $\mathcal{P}\to\mathcal{T}_{S}$. Our goal is to show that there exists a base change $\tilde{\mathcal{T}_{S}}\to\mathcal{T}_{S}$ such that $\mathcal{P}\times_{\mathcal{T}_{S}}\tilde{\mathcal{T}_{S}}$ splits as $\mu$ copies of $\tilde{\mathcal{T}_{S}}$. As a first step, we consider the following (somewhat absurd non-sensical) base change. ${\mathcal{P}\times_{\mathcal{T}_{S}}\mathcal{P}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{P}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{P}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{T}_{S}}$ The diagonal $\Delta$ provides a section of the degree $\mu$ map $\mathcal{P}\times_{\mathcal{T}_{S}}\mathcal{P}\to\mathcal{P}$, hence $\mathcal{P}\times_{\mathcal{T}_{S}}\mathcal{P}$ splits as $\Delta+\mathcal{P}^{\prime}$ where $\mathcal{P}^{\prime}$ comes with a degree $\mu-1$ map to $\mathcal{P}$. Then, by induction on $\mu$, one may construct a finite cover $\tilde{\mathcal{P}}\to\mathcal{P}$ such that $\mathcal{P}^{\prime}\times_{\mathcal{P}}\tilde{\mathcal{P}}$ splits as $\mu-1$ copies of $\tilde{\mathcal{P}}$, hence $\mathcal{P}\times_{\mathcal{T}_{S}}\tilde{\mathcal{P}}$ splits as $\mu$ copies of $\tilde{\mathcal{P}}$. Thus $\tilde{\mathcal{P}}$ is suitable as the $\tilde{\mathcal{T}_{S}}$ we were looking for. We point out the (non-accidental) similarity of this construction with that of a splitting field for a given polynomial $P$ in one indeterminate $X$. The first base change in our construction is the analogue of the field extension $\mathbf{k}[X]/(P)$. Also note that, in general, the map $\tilde{\mathcal{T}_{S}}\to\mathcal{T}_{S}$ will have degree $\mu!$. (in other words the factors $\Phi^{i}_{p^{\prime},p^{\prime\prime}}$ have their coefficients in a finite extension of $\mathbf{k}[p^{\prime},p^{\prime\prime}]$). Then, by the multiplicativity property, see [15, § 5.7, p. 154], the resultant $R$ splits as the product of the $\mu$ resultants $R_{i}=\mathrm{Res}(T_{p^{\prime}},\Phi^{i}_{p^{\prime},p^{\prime\prime}},L^{\prime},L^{\prime\prime})$. By Lemma (3.3), for all $i$ the resultant $R_{i}$ vanishes wherever $D_{L^{\prime},L^{\prime\prime}}$ vanishes on $\mathcal{T}_{S}$, or rather on $\tilde{\mathcal{T}_{S}}$, as both forms $T_{p^{\prime}}$ and $\Phi^{i}_{p^{\prime},p^{\prime\prime}}$ vanish along the line $\langle p^{\prime},p^{\prime\prime}\rangle$. By Lemma (3.4), this implies that $D_{L^{\prime},L^{\prime\prime}}$ divides $R_{i}$ as a function of $(p^{\prime},p^{\prime\prime})\in\tilde{\mathcal{T}_{S}}$. Eventually, the upshot is that $(D_{L^{\prime},L^{\prime\prime}})^{\mu}$ divides $R$, and since both $D_{L^{\prime},L^{\prime\prime}}$ and $R$ are well-defined as functions of $(p^{\prime},p^{\prime\prime})\in{\mathcal{T}_{S}}$, so is the quotient. Therefore, there exists a multi-homogeneous polynomial $R^{\prime}$ in the sets of variables $(L^{\prime},L^{\prime\prime},p^{\prime},p^{\prime\prime})$ such that after restriction to $\check{\mathbf{P}}^{3}\times\check{\mathbf{P}}^{3}\times\mathcal{T}_{S}$ one has $R=(D_{L^{\prime},L^{\prime\prime}})^{\mu}R^{\prime}$. Computing degrees, one finds that $R^{\prime}$ has multi-degree $(0,0,\mu(d-2)+\lambda,0)$ in $(L^{\prime},L^{\prime\prime},p^{\prime},p^{\prime\prime})$, which proves the theorem. ∎
Study of a pencil of planes orthogonal to a tangent line
(3.1)General setup
We recall that $\mathbf{k}$ is assumed to be an algebraically closed field of characteristic 0. We let $S\subseteq\mathbf{P}^{3}$ be a smooth surface of degree $d$, defined by a homogeneous polynomial $F(X,Y,Z,W)$. We consider generic $4$-tuples $\hat{p}^{\prime}=(X^{\prime},Y^{\prime},Z^{\prime},W^{\prime})$, $\hat{p}^{\prime\prime}=(X^{\prime\prime},Y^{\prime\prime},Z^{\prime\prime},W^{\prime\prime})$, $\hat{p}=(X,Y,Z,W)$, and call $p^{\prime},p^{\prime\prime},p$ the corresponding points in $\mathbf{P}^{3}$ (beware the unusual distribution of the prime decorations). The choice of $\hat{p}^{\prime},\hat{p}^{\prime\prime},\hat{p}$ defines a system of homogeneous coordinates $(\alpha:\beta:\gamma)$ on the plane $\langle p^{\prime},p^{\prime\prime},p\rangle$ generated by $^{\prime}, p′′p^{\prime\prime}$ and $p$. We consider
[…]
$=\alpha^{d}\cdot F(\hat{p}^{\prime})+\alpha^{d-1}\cdot\mathrm{D}^{\beta\hat{p}^{\prime\prime}+\gamma\hat{p}}F(\hat{p}^{\prime})+\frac{1}{2}\alpha^{d-2}\cdot\mathrm{D}^{(\beta\hat{p}^{\prime\prime}+\gamma\hat{p})^{2}}F(\hat{p}^{\prime})\ \ \textrm{mod}\ (\beta,\gamma)^{3}$
$=\alpha^{d}\cdot F(\hat{p}^{\prime})+\alpha^{d-1}\bigl(\beta\mathrm{D}^{\hat{p}^{\prime\prime}}F(\hat{p}^{\prime})+\gamma\mathrm{D}^{\hat{p}}F(\hat{p}^{\prime})\bigr)$
$\qquad+\frac{1}{2}\alpha^{d-2}\bigl(\beta^{2}\mathrm{D}^{\hat{p}^{\prime\prime 2}}F(\hat{p}^{\prime})+2\beta\gamma\mathrm{D}^{\hat{p}^{\prime\prime}\hat{p}}F(\hat{p}^{\prime})+\gamma^{2}\mathrm{D}^{\hat{p}^{2}}F(\hat{p}^{\prime})\bigr)\quad\ \textrm{mod}\ (\beta,\gamma)^{3}.$
Considered as a homogeneous polynomial in the variables $(\alpha,\beta,\gamma)$, this is the equation of the hyperplane section of $S$ by $\langle p^{\prime},p^{\prime\prime},p\rangle$.
(3.2)Degree of the dual surface
As a warm-up, let us derive the degree of the dual surface $S^{\vee}\subseteq\smash{\check{\mathbf{P}}^{3}}$ of $S$ from the relevant homogeneity property of the discriminant. We specialize $p^{\prime\prime},p$ to two general points of $\mathbf{P}^{3}_{\mathbf{k}}$. Then the coefficients of $f(\alpha,\beta,\gamma)$ are polynomials in the set of indeterminates $\hat{p}^{\prime}=(X^{\prime},Y^{\prime},Z^{\prime},W^{\prime})$, and so is the discriminant $\mathrm{Disc}(f)$ with respect to the variables $\alpha,\beta,\gamma$. The latter vanishes at a point $p^{\prime}\in\mathbf{P}^{3}_{\mathbf{k}}$ if and only if the plane $\langle p^{\prime},p^{\prime\prime},p\rangle$ is tangent to $S$. Thus the zero locus $\mathrm{V}\bigl(\mathrm{Disc}(f)\bigr)\subseteq\mathbf{P}^{3}_{\mathbf{k}}$ is the union of all planes tangent to $S$ and containing the line $\langle p^{\prime\prime},p\rangle$; in particular the degree of $\mathrm{Disc}(f)$ (as a polynomial in $p^{\prime}$, what else?) is the number of tangent planes to $S$ containing the line $\langle p^{\prime\prime},p\rangle$, which is the degree of the dual surface $S^{\vee}\subseteq\smash{\check{\mathbf{P}}^{3}}$. This is also the number of points in the intersection of $S^{\vee}$ with the line $(p^{\prime\prime},p)^{\perp}\subseteq\smash{\check{\mathbf{P}}^{3}}$, the latter parametrizing planes passing through $p^{\prime\prime}$ and $p$.
The coefficient of $f(\alpha,\beta,\gamma)$ in $\alpha^{a}\beta^{b}\gamma^{c}$ is a polynomial of degree $a$ in the set of indeterminates $\hat{p}^{\prime}$, as can be seen on the Taylor–Newton expansion, the first terms of which are given in Display (3.1.1). Therefore, by (2.6.1), the discriminant $\mathrm{Disc}(f)$ is homogeneous of degree $d(d-1)^{2}$ in $\hat{p}^{\prime}$, and the dual surface $S^{\vee}$ has degree $d^{\vee}=d(d-1)^{2}$ in $\smash{\check{\mathbf{P}}^{3}}$.
(3.3) Remark.
If we let all three points (or $4$-tuples) $p^{\prime},p^{\prime\prime},p$ be generic, then $\mathrm{Disc}(f(\alpha,\beta,\gamma))$ is tri-homogeneous with respect to the three sets of indeterminates $p^{\prime},p^{\prime\prime},p$, of tri-degree
$(d^{\vee},d^{\vee},d^{\vee})=(d(d-1)^{2},d(d-1)^{2},d(d-1)^{2}),$
as follows from the above analysis. Note that obviously $p^{\prime},p^{\prime\prime},p$ play symmetric roles in the definition of $f$.
In order to enumerate bitangent planes, we will take points $p^{\prime},p^{\prime\prime}$ such that the line $\langle p^{\prime},p^{\prime\prime}\rangle$ is tangent to $S$ at $p^{\prime}$, and consider a reduced discriminant of the corresponding $f(\alpha,\beta,\gamma)$. Thus we need to specialize our setup a little.
(3.4)The “tangent line” setup
In the situation set-up in Paragraph (3.1) above, we specialize $p^{\prime}$ to the general point of $S$, and $p^{\prime\prime}$ to the general point of $\mathbf{T}_{p^{\prime}}S$. Then $F(\hat{p}^{\prime})=\mathrm{D}^{\hat{p}^{\prime\prime}}F(\hat{p}^{\prime})=0$, so (3.1.1) reduces to
(3.4.1) $f(\alpha,\beta,\gamma)=T\alpha^{d-1}\gamma+\frac{1}{2}\alpha^{d-2}(A\beta^{2}+2B\beta\gamma+C\gamma^{2})\quad\ \textrm{mod}\ (\beta,\gamma)^{3},$
where $T=\mathrm{D}_{\hat{p}}f(\hat{p}^{\prime})$, $A=\mathrm{D}_{\hat{p}^{\prime\prime 2}}f(\hat{p}^{\prime})$, $B=\mathrm{D}_{\hat{p}^{\prime\prime}\hat{p}}f(\hat{p}^{\prime})$, $C=\mathrm{D}_{\hat{p}^{2}}f(\hat{p}^{\prime})$. Note in particular that, as homogeneous polynomials in the set of indeterminates $\hat{p}=(X,Y,Z,W)$, $T$ and $C$ are the equations of the tangent plane $\mathbf{T}_{p^{\prime}}S$, and the polar quadric $\mathrm{D}^{2}S(p^{\prime})$ of $S$ at $p^{\prime}$, respectively.
We consider the discriminant of the plane curve $S\cap\langle p^{\prime},p^{\prime\prime},p\rangle$, i.e., the discriminant of $f\in A[\alpha,\beta,\gamma]$ with $A=\mathbf{k}[X,Y,Z,W]$; by Formula (a), or rather Formula (2.1.2), it writes
(3.4.2) $\mathrm{Disc}(f)=-\frac{1}{2}\,T^{2}\bigl(A(B^{2}-AC)^{2}\,\Phi+T\,\Psi\bigr),\quad\text{\text{with}}\quad\Phi=\operatorname{redDisc}_{d}^{2}(f).$
As explained in the previous paragraphs, it vanishes if and only if the plane $\langle p^{\prime},p^{\prime\prime},p\rangle$ is tangent to $S$; It is a trihomogeneous polynomial in the variables $\hat{p}^{\prime}$, $\hat{p}^{\prime\prime}$ and $\hat{p}$; as a polynomial in $p$, it vanishes along the $d^{\vee}$ planes (counted with multiplicities) tangent to $S$ in the pencil $(p^{\prime},p^{\prime\prime})^{\perp}\subseteq\smash{\check{\mathbf{P}}^{3}}$ of planes containing the line $\langle p^{\prime},p^{\prime\prime}\rangle\subseteq\mathbf{P}^{3}$.
(3.5) Remark.
The fact that $T^{2}$ factors out of $\mathrm{Disc}(f)$ gives an algebraic proof of the fact that the tangent plane $\mathbf{T}_{p^{\prime}}S$ appears with multiplicity $\geqslant 2$ in the scheme $(p^{\prime},p^{\prime\prime})^{\perp}\cap S^{\vee}$; or rather, the point $(\mathbf{T}_{p^{\prime}}S)^{\perp}\in\smash{\check{\mathbf{P}}^{3}}$ appears with multiplicity $\geqslant 2$ in the scheme $(p^{\prime},p^{\prime\prime})^{\perp}\cap S^{\vee}$.
In other words, if a line $\Lambda\subseteq\mathbf{P}^{3}$ is tangent to $S$ at some point $p$, then the line $\Lambda^{\perp}\subseteq\smash{\check{\mathbf{P}}^{3}}$ is tangent to the dual surface $S^{\vee}$ at the point $(\mathbf{T}_{p}S)^{\perp}$.
Salmon’s idea is that if the tangent plane $\mathbf{T}_{p^{\prime}}S$ is also tangent to $S$ at some point different from $p^{\prime}$, then it should appear with multiplicity >2>2 in $(p^{\prime},p^{\prime\prime})^{\perp}\cap S^{\vee}$; equivalently, $\mathrm{Disc}(f)$ should be divisible by $T^{3}$ in the above setup.
Therefore, we shall now discuss the various possibilities for $T$ to divide $A(B^{2}-AC)^{2}\,\Phi$, in the notation of Equation (3.4.2).
(3.6)Vanishing of $A$ and biduality for inflectional tangents
The polynomial $A$ has degree 0 in $p$, therefore it is divisible by $T$ if and only if it is identically zero. By definition $A=\mathrm{D}^{\hat{p}^{\prime\prime 2}}f(\hat{p}^{\prime})$, so its vanishing is equivalent to the point $p^{\prime\prime}$ being on the polar quadric $\mathrm{D}_{{p^{\prime}}^{2}}S$ of $S$ at $p^{\prime}$. Since $p^{\prime\prime}\in\mathbf{T}_{p^{\prime}}S$, this in turn is equivalent to the line $\langle p^{\prime},p^{\prime\prime}\rangle$ being one of the two inflectional tangents of $S$ at $p^{\prime}$ (i.e., the two lines $\Lambda$ having intersection multiplicity with $S$ at $p^{\prime}$ at least $3$; in symbols, $i(\langle p^{\prime},p^{\prime\prime}\rangle,S)_{p^{\prime}}\geqslant 3$).
We thus find an algebraic proof of the fact that
$\text{if}\quad i\bigl(\langle p^{\prime},p^{\prime\prime}\rangle,S\bigr)_{p^{\prime}}\geqslant 3,\quad\text{then}\quad i\bigl((p^{\prime},p^{\prime\prime})^{\perp},S^{\vee}\bigr)_{(\mathbf{T}_{p^{\prime}}S)^{\perp}}\geqslant 3,$
which is a manifestation of biduality. Geometrically, if a line $\Lambda\subseteq\mathbf{P}^{3}$ is an inflectional tangent to $S$ at $p^{\prime}$ (i.e., $L$ is the tangent line to one of the two local branches of $\mathbf{T}_{p}S\cap S$ at $p^{\prime}$), then its orthogonal $\Lambda^{\perp}\subseteq\check{\mathbf{P}}^{3}$ is an inflectional tangent to $S^{\vee}$ at the point $(\mathbf{T}_{p^{\prime}}S)^{\perp}$.
As a side remark note that when $A=0$, all the curves cut out on $S$ by a member of the pencil $(p^{\prime},p^{\prime\prime})^{\perp}$ have an inflection point at $p^{\prime}$.
We shall analyze the divisibilities of $B^{2}-AC$ and $\Phi$ by $T$ using Theorem (3.2). To see that the latter result indeed applies, we note that for $(p^{\prime},p^{\prime\prime})\in\mathcal{T}_{S}$ (see the notation in Paragraph (3.1)), the discriminant in Display (3.4.2), $T^{2}A(B^{2}-AC)^{2}\Phi+T^{3}\Psi$, defines as a homogeneous polynomial in the variable $p$ a hypersurface consisting of $d^{\vee}$ planes, counted with multiplicities, all containing the line $\langle p^{\prime},p^{\prime\prime}\rangle$. This implies that so does its homogeneous piece of lowest degree with respect to any grading, as the piece of lowest degree of a product is the product of the pieces of lowest degree. For the Zariski grading, see subsection 2.3, the piece of lowest degree of the discriminant is $T^{2}A(B^{2}-AC)^{2}\Phi$. The upshot is that $T^{2}A(B^{2}-AC)^{2}\Phi$, as a polynomial in $p$, defines a sum of planes, all containing the line $\langle p^{\prime},p^{\prime\prime}\rangle$, which implies that so do all the factors $T,A,B^{2}-AC,\Phi$.
Note that this is obvious for $T$, which defines the plane $\mathbf{T}_{p^{\prime}}S$. The polynomial $A$, on the other hand, is independent on the variable $p$, hence defines either the whole space, or the empty set.
(3.8)Divisibility of $B^{2}-AC$ by $T$ and biduality at parabolic points
The polynomial $B^{2}-AC$ is tri-homogeneous of tri-degree (2(d2),2,2)(2(d-2),2,2) in the variables $(p^{\prime},p^{\prime\prime},p)$. By Theorem (3.2), arguing as in Paragraph (3.7) above, there exists a homogeneous polynomial $H$ of degree 4(d2)4(d-2) in the variable $p^{\prime}$, with constant coefficients, such that for fixed $p^{\prime}\in S$ and $p^{\prime\prime}\in\mathbf{T}_{p^{\prime}}S$, $T$ divides $B^{2}-AC$ as polynomials in $p$ if and only if $H(p^{\prime})=0$.
The polynomial $H$ is, in fact, the Hessian determinant. Indeed, it follows from (3.4.1) that the tangent cone of the section of $S$ by its tangent hyperplane at $p^{\prime}$, $S\cap\mathbf{T}_{p^{\prime}}S$, is defined by the quadratic form given by the symmetric matrix
$\begin{pmatrix}A&B\\ B&C\end{pmatrix}$.
Thus, $B^{2}-AC$ is zero modulo $T$ if and only if the curve $S\cap\mathbf{T}_{p^{\prime}}S$ has a degenerate tangent cone at $p^{\prime}$, i.e., $p^{\prime}$ is a parabolic point of $S$.
Geometrically this is explained as follows. If $p^{\prime}$ is a parabolic point of $S$, then the point $(\mathbf{T}_{p^{\prime}}S)^{\perp}\in\check{\mathbf{P}}^{3}$ sits on the cuspidal double curve of $S^{\vee}$, and for $p^{\prime\prime}\in\mathbf{T}_{p^{\prime}}S-\{p^{\prime}\}$, the line $(p^{\prime},p^{\prime\prime})^{\perp}\subseteq\check{\mathbf{P}}^{3}$ is contained in the tangent cone of $S^{\vee}$ at the point $(\mathbf{T}_{p^{\prime}}S)^{\perp}\in\check{\mathbf{P}}^{3}$, hence $i((p^{\prime},p^{\prime\prime})^{\perp},S^{\vee})_{(\mathbf{T}_{p^{\prime}}S)^{\perp}}\geqslant 3$.
(3.9)Divisibility of the reduced discriminant $\Phi$ by $T$
The discriminant in Display (3.4.2) is tri-homogeneous in the variables $(p^{\prime},p^{\prime\prime},p)$ with respect to the standard grading, of degree $d^{\vee}=d(d-1)^{2}$ with respect to all three variables, as we have observed above. This implies that also the homogeneous piece of (3.4.2) of lowest degree with respect to the Zariski grading is tri-homogeneous of tri-degree $(d^{\vee},d^{\vee},d^{\vee})$ for the standard grading. Moreover, $T$, AA, and $B^{2}-AC$ have respective tri-degrees (d1,0,1)(d-1,0,1), (d2,2,0)(d-2,2,0), and (2(d2),2,2)(2(d-2),2,2) in $(p^{\prime},p^{\prime\prime},p)$. Thus, computing degrees, one finds that $\Phi$ has tri-degree $(\lambda,\mu,\mu)$, with
$\lambda=(d-2)(d^{2}-6)\quad\text{and}\quad\mu=d^{3}-2d^{2}+d-6.$
Since $\Phi$ is $\operatorname{redDisc}_{d}^{2}(f)$ with respect to the indeterminate $\alpha$, and the coefficient of $f(\alpha,\beta,\gamma)$ in $\alpha^{a}\beta^{b}\gamma^{c}$ is a polynomial of degree $a$ in the indeterminates $p^{\prime}$, the fact that $\Phi$ is homogeneous in $p^{\prime}$ and its degree follow directly from (2.14.1); to wit, in our situation the degree in (2.14.1) reads
$d(d-1)^{2}-\bigl[7d-12\bigr]=d^{3}-2d^{2}-6d+12=(d-2)(d^{2}-6).$
The homogeneity and degrees in $p$ and $p^{\prime\prime}$ however do not directly follow from Proposition (2.14), as for instance the coefficients in $A\beta^{2}+2B\beta\gamma+C\gamma^{2}$ do not all have the same weights in $p$ or $p^{\prime\prime}$. It is noticeable that $T^{2}A(B^{2}-AC)^{2}$ is homogeneous in $p$ and $p^{\prime\prime}$ because the weights of $T^{2}$ and $A$ balance each other; this is a very elementary reason why $T$ “should” appear with a square in Formula (3.4.2). The degree of $\operatorname{redDisc}_{d}^{2}(f)$ with the weights as in (2.14.2) is
$2\bigl[d(d-1)^{2}-6\bigr]=2\bigl[d^{3}-2d^{2}+d-6\bigr],$
and one observes that its degrees in $p$ and $p^{\prime\prime}$ respectively are one half of this, thus this degree “distributes equally” between $p$ and $p^{\prime\prime}$.
In conclusion, it follows as in Paragraph (3.8) from Theorem (3.2) that there exists a homogeneous polynomial $K$ of degree $(d-2)(d^{3}-d^{2}+d-12)$ in the variable $p^{\prime}$, with constant coefficients, such that for fixed $p^{\prime}\in S$ and $p^{\prime\prime}\in\mathbf{T}_{p^{\prime}}S$, $T$ divides $\Phi$ if and only if $K(p^{\prime})=0$.
We have thus arrived at the following statement.
(3.10) Theorem.
Let $S$ be a smooth, degree $d$, surface in $\mathbf{P}^{3}$. There is a hypersurface $V(K)$ of degree $(d-2)(d^{3}-d^{2}+d-12)$, the intersection of which with $S$ is the locus of tangency points of planes bitangent to $S$.
It has to be recognized that the proof given above of the latter statement rests on the not fully justified fact that $T$ divides $\Phi=\operatorname{redDisc}(f)$ if and only if the tangent plane $\mathbf{T}_{p^{\prime}}S$ is also tangent to $S$ at some additional point. The “if” part is established by [II, Theorem **], which is a result due to Dimca [9], but what is really needed is the “only if” part. Evidence in favour of that assertion is provided by the irreducibility of the reduced discriminant $\Phi$ (which, as already mentioned, follows from the fact that the reduced discriminant is a toric discriminant; see also [4]).
(3.11) Corollary.
The ordinary double curve of the dual surface $S^{\vee}$ has degree $\frac{1}{2}d(d-1)(d-2)(d^{3}-d^{2}+d-12).$
Proof.
Let $p^{\prime\prime}\in\mathbf{P}^{3}$ be a general point. The locus of those points $p^{\prime}\in S$ such that there exists a plane through $p^{\prime\prime}$ tangent to $S$ at $p^{\prime}$ is the apparent boundary $\mathrm{D}^{p^{\prime\prime}}S\cap S$. Therefore, by Theorem (3.10), the locus of points $p^{\prime}\in S$ such that $\mathbf{T}_{p^{\prime}}S$ is bitangent and passes through $p^{\prime\prime}$ is $\mathrm{D}_{p^{\prime\prime}}S\cap S\cap V(K)$. Now for each bitangent plane there are two tangency points $p^{\prime}$, so the number of bitangent planes passing through $p^{\prime\prime}$ is
$\frac{1}{2}\cdot\deg(S)\cdot\deg(\mathrm{D}^{p^{\prime\prime}}S)\cdot\deg K.$
Generalization to hypersurfaces of arbitrary dimension
In fact, Salmon’s procedure works in arbitrary dimension, using the generalization (2.2.6) of formula (a). The arguments are direct generalizations, so we are going to be sketchy.
Let $V=\mathrm{V}(F)$ be a smooth hypersurface of degree $d$ in $\mathbf{P}^{n}$, and consider $n$ points $p_{1},\ldots,p_{n-1},p$ (or rather tuples $\hat{p}_{1},\ldots,\hat{p}_{n-1},\hat{p}\in\mathbf{k}^{n+1}$). The points $p_{1},\ldots,p_{n-1}$ (if in linear general position, i.e., if $\hat{p}_{1},\ldots,\hat{p}_{n-1}$ are linearly independent) define a pencil of hyperplanes in $\mathbf{P}^{n}$, and for all $p$ (not in $\langle p_{1},\ldots,p_{n-1}\rangle$), the tuples $\hat{p}_{1},\ldots,\hat{p}_{n-1},\hat{p}$ define a system of homogeneous coordinates $(\alpha_{1}:\ldots:\alpha_{n-1}:\alpha)$ on the member $H_{p}$ of the pencil determined by $p$, i.e., $\langle p_{1},\ldots,p_{n-1},p\rangle$. In this system of coordinates, the hyperplane section $H_{p}\cap V$ is the zero locus of the polynomial
[…]
$=\sum_{k=0}^{d}\,\frac{1}{k!}\,\alpha_{1}^{d-k}\,\mathrm{D}^{(\alpha_{2}\hat{p}_{2}+\cdots+\alpha\hat{p})^{k}}F(\hat{p}_{1}).$
We shall assume that $p_{1}\in V$ and $p_{2},\ldots,p_{n-1}\in\mathbf{T}_{p_{1}}V$, equivalently
$F(p_{1})=\mathrm{D}^{p_{2}}F(p_{1})=\cdots=\mathrm{D}^{p_{n-1}}F(p_{1})=0.$
Then,
$f(\alpha_{1},\ldots,\alpha_{n-1},\alpha)=T\alpha_{1}^{d-1}\alpha+\frac{1}{2}\alpha_{1}^{d-2}f_{2}(\alpha_{2},\ldots,\alpha_{n},\alpha)\quad\ \textrm{mod}\ (\alpha_{2},\ldots,\alpha)^{3}$
where $T=\mathrm{D}_{\hat{p}}f(\hat{p}_{1})$ and $f_{2}(\alpha_{2},\ldots,\alpha_{n},\alpha)=\mathrm{D}^{(\alpha_{2}\hat{p}_{2}+\cdots+\alpha\hat{p})^{2}}F(\hat{p}_{1})$. By Formula (2.2.6), the generalization of (a), one has
$\mathrm{Disc}(f)=T^{2}\cdot\mathrm{Disc}(\bar{f}_{2})\cdot\mathrm{Disc}(f_{2})^{2}\cdot\Phi\quad\ \textrm{mod}\ T^{3},$
with $\bar{f}_{2}(\alpha_{2},\ldots,\alpha_{n})=f_{2}(\alpha_{2},\ldots,\alpha_{n},0)$ and $\Phi=\operatorname{redDisc}_{d}^{2}(f)$. Our task is to analyze the divisibilities of $\mathrm{Disc}(f_{2})$ and $\Phi$ by $T$ ($\bar{f}_{2}$ has degree 0 in $p$, hence $T$ divides $\mathrm{Disc}(\bar{f}_{2})$ if and only if $\mathrm{Disc}(\bar{f}_{2})=0$).
Theorem (3.2) (Salmon’s elimination trick) generalizes as follows. Given n1n-1 linear forms $L_{1},\ldots,L_{n-1}\in\mathbf{P}^{n}$, the line $\mathrm{V}(L_{1},\ldots,L_{n-1})$ intersects the (n2)(n-2)-dimensional linear space $\langle p_{1},\ldots,p_{n-1}\rangle$ if and only if
$D_{L_{1},\ldots,L_{n-1}}(p_{1},\ldots,p_{n-1})\overset{\text{def.}}{=}\det\bigl(L_{i}(p_{j})\bigr)_{1\leqslant i,j\leqslant n-1}$
vanishes.
Let
$\mathcal{T}_{V}:=\{(p_{1},\ldots,p_{n-1})\in(\mathbf{P}^{n})^{n-1}:p_{1}\in V\text{ and }p_{2},\ldots,p_{n-1}\in\mathbf{T}_{p_{1}}V\},$
and consider a multihomogeneous polynomial $\Phi_{p_{1},\ldots,p_{n-1}}(p)$ of multidegree $(\lambda,\mu,\ldots,\mu)$ in the sets of variables $p_{1},\ldots,p_{n-1},p$, such that for all $(p_{1},\ldots,p_{n-1})\in\mathcal{T}_{V}$, the hypersurface $\mathrm{V}(\Phi_{p_{1},\ldots,p_{n-1}})$ consists of $\mu$ hyperplanes, counted with multiplicities, all containing the (n2)(n-2)-plane $\langle p_{1},\ldots,p_{n-1}\rangle$.
Then, the locus of those $p_{1}\in V$ such that there exist $p_{2},\ldots,p_{n-1}\in\mathbf{T}_{p_{1}}V$ such that $\mathbf{T}_{p_{1}}X$ is a component of $\mathrm{V}(\Phi_{p_{1},\ldots,p_{n-1}})$ is cut out on $V$ by a hypersurface of degree $\lambda+\mu(d-2)$.
One may then argue as in (3.7) to analyze the divisibilities of $\mathrm{Disc}(f_{2})$ and $\Phi$ by $T$.
For $\mathrm{Disc}(f_{2})$, we find as before that it is divisible by $T$ if and only if $p_{1}$ lies on the Hessian hypersurface of $V$, since $f_{2}\pmod{T}$ defines the tangent cone at $p_{1}$ of the tangential hyperplane section $\mathbf{T}_{p_{1}}V\cap V$, hence $\mathrm{Disc}(f_{2})$ vanishes modulo $T$ if and only if the latter tangent cone is singular, i.e., if and only if $p_{1}$ is a parabolic point of $V$.
And indeed $\mathrm{Disc}(f_{2})$, being the determinant of the symmetric matrix
$\begin{pmatrix}\mathrm{D}^{(p_{2})^{2}}F(p_{1})&\cdots&\mathrm{D}^{p_{2}p_{n-1}}F(p_{1})&\mathrm{D}^{p_{2}p}F(p_{1})\\ \vdots&\ddots&\vdots&\vdots\\ \mathrm{D}^{p_{n-1}p_{2}}F(p_{1})&\cdots&\mathrm{D}^{(p_{n-1})^{2}}F(p_{1})&\mathrm{D}^{p_{n-1}p}F(p_{1})\\ \mathrm{D}^{pp_{2}}F(p_{1})&\cdots&\mathrm{D}^{pp_{n-1}}F(p_{1})&\mathrm{D}^{p^{2}}F(p_{1})\end{pmatrix}$,
has multidegree $\bigl((n-1)(d-2),2,\ldots,2\bigr)$ in $p_{1},\ldots,p_{n-1},p$, so the argument of (3.2) produces a homogeneous polynomial $H$ of degree
(n1)(d2)+2(d2)=(n+1)(d2)(n-1)(d-2)+2(d-2)=(n+1)(d-2)
in the variable $p_{1}$.
The analysis of the divisibility of the reduced discriminant $\Phi$ by $T$, on the other hand, gives rise to the couple-nodal polynomial $K$.
Let us first compute the multidegree of $\Phi$ in the variables $p_{1},\ldots,p_{n-1},p$. This goes as in Paragraph (3.9). First of all, $\mathrm{Disc}(f)$ is $n$-homogeneous of $n$-degree $(d^{\vee},\ldots,d^{\vee})$ in $p_{1},\ldots,p_{n-1},p$, with $d^{\vee}=d(d-1)^{n-1}$, hence so is its homogeneous piece of lowest degree with respect to the Zariski grading
$T^{2}\,\mathrm{Disc}(\bar{f}_{2})\,\mathrm{Disc}(f_{2})^{2}\,\Phi.$
On the other hand, $T$ has $n$-degree $(d-1,0,\ldots,0,1)$, while $\mathrm{Disc}(f_{2})$ has $n$-degree $((n-1)(d-2),2,\ldots,2)$ as we saw in Paragraph (3.3) above. The same computation gives the $n$-degree of $\mathrm{Disc}(\bar{f}_{2})$, namely $((n-2)(d-2),2,\ldots,2,0)$. Eventually, one finds that $\Phi$ has degrees
$d\bigl[(d-1)^{n-1}-1\bigr]-3(n-1)(d-2)\quad\text{and}\quad d(d-1)^{n-1}-6$
in $p_{1}$, and $p_{2},\ldots,p_{n-1},p$, respectively (note that the former degree is divisible by d2d-2). As in Paragraph (3.9), the degree in $p_{1}$ of $\operatorname{redDisc}_{d}^{2}(f)$ is given by (2.14.1), while the degree with respect to the grading of (2.14.2) is equally distributed between the degrees in $p_{2},\ldots,p_{n-1},p$ respectively.
Eventually, by the result in Paragraph (3.2), there exists a polynomial $K$ in $p_{1}$, homogeneous of degree
$d\bigl[(d-1)^{n}-1\bigr]-3(n+1)(d-2)=(d-2)\Bigl(d\cdot\frac{(d-1)^{n}-1}{d-2}-3(n+1)\Bigr),$
such that $T$ divides $\Phi=\operatorname{redDisc}_{d}^{2}(f)$ if and only if $p_{1}\in V$ lies on the hypersurface $\mathrm{V}(K)$. We call $K$ the couple-nodal polynomial. One thus obtains the following result.
(3.5) Theorem.
Let $V$ be a smooth degree $d$ hypersurface in $\mathbf{P}^{n}$, $n>1$. The number of hyperplanes bitangent to $V$ passing through n2n-2 fixed general points in $\mathbf{P}^{n}$ is
$\frac{1}{2}d(d-1)^{n-2}(d-2)\Bigl(d\cdot\frac{(d-1)^{n}-1}{d-2}-3(n+1)\Bigr).$
For $n=3$ one recovers Theorem (3.10), and for $n=2$ the number of bitangents to a smooth plane curve of degree $d$, viz.
$\frac{1}{2}(d^{\vee}-1)(d^{\vee}-2)-3d(d-2)-\frac{1}{2}(d-1)(d-2)=\frac{1}{2}d(d-2)(d-3)(d+3).$
Number of bitangent lines and generalizations
In this section, we elaborate on the computation by Salmon of the number of bitangent lines to a surface in $\mathbf{P}^{3}$ passing through a general point. We give his proof in Section 4.1, and a generalization to hypersurfaces in $\mathbf{P}^{n}$ in the next Section 4.2. These proofs only involve the reduced discriminant of polynomials in two indeterminates, which is easily expressed as a plain discriminant, see Example (2.10). We give however another generalization in Section 4.3, for which it is necessary to consider reduced discriminants of polynomials with an arbitrary number of indeterminates. The two statements of Sections 4.1 and 4.2 are instances of a vast generalization by Fehér and Juhász [10], which is described in [III, Theorem **].
Number of bitangent lines
The Grassmannian of lines in $\mathbf{P}^{3}$ has dimension $4$. Passing through a fixed point imposes $2$ conditions to a line in $\mathbf{P}^{3}$, and being tangent to a surface (at an unprescribed point) imposes $1$ condition, so one expects finitely many bitangent lines to a surface passing through a general point in $\mathbf{P}^{3}$. In this subsection we prove the following, along the lines of [17, §279].
(4.1) Theorem.
Let $S$ be a smooth surface of degree $d$ in $\mathbf{P}^{3}$, and $p\in\mathbf{P}^{3}$ a general point. The number of lines bitangent to $S$ and passing through $p$ is
(4.1.1) $\frac{1}{2}d(d-1)(d-2)(d-3).$
Salmon’s strategy is similar in spirit to that exposed in subsection 3.2. The fundamental fact is the following. Consider as before
$\mathcal{T}_{S}=\bigl\{(p^{\prime},p^{\prime\prime})\in\mathbf{P}^{3}\times\mathbf{P}^{3}:p^{\prime}\in S\text{ and }p^{\prime\prime}\in\mathbf{T}_{p^{\prime}}S-\{p^{\prime}\}\bigr\};$
equivalently, $(p^{\prime},p^{\prime\prime})$ is in $\mathcal{T}_{S}$ if and only if $\langle p^{\prime},p^{\prime\prime}\rangle$ is a line, tangent to $S$ at $p^{\prime}$.
(4.2) Proposition.
There exists a polynomial $R$, bihomogeneous in $p^{\prime},p^{\prime\prime}$ with respective degrees (d2)(d3)(d-2)(d-3) and (d+2)(d3)(d+2)(d-3), such that the locus of points $(p^{\prime},p^{\prime\prime})\in\mathcal{T}_{S}$ for which the line $\langle p^{\prime},p^{\prime\prime}\rangle$ is bitangent to $S$ is cut out on $\mathcal{T}_{S}$ by $V(R)$.
Proof.
Let $F$ be an equation of $S\subseteq\mathbf{P}^{3}$. For $(p^{\prime},p^{\prime\prime})\in\mathcal{T}_{S}$, we consider the homogeneous polynomial in $(\alpha,\beta)$
$f(\alpha,\beta)$ $=F(\alpha p^{\prime}+\beta p^{\prime\prime})$
$=\alpha^{d}F(p^{\prime})+\alpha^{d-1}\beta\mathrm{D}^{p^{\prime\prime}}F(p^{\prime})+\frac{1}{2}\alpha^{d-2}\beta^{2}\mathrm{D}^{(p^{\prime\prime})^{2}}F(p^{\prime})+\cdots+\beta^{d}F(p^{\prime\prime})$
(4.2.1) $=\beta^{2}\Bigl(\frac{1}{2}\alpha^{d-2}\mathrm{D}^{(p^{\prime\prime})^{2}}F(p^{\prime})+\cdots+\beta^{d-2}F(p^{\prime\prime})\Bigr)$
(this is an abuse of notation: actually one should consider two liftings $\hat{p}^{\prime},\hat{p}^{\prime\prime}\in\mathbf{k}^{4}$ of $p^{\prime}$ and $p^{\prime\prime}$ respectively, and $F(\alpha\hat{p}^{\prime}+\beta\hat{p}^{\prime\prime})$); the last equality comes from the fact that $\langle p^{\prime},p^{\prime\prime}\rangle$ is tangent to $S$ at $p^{\prime}$. Let $\vphantom{f}{}^{\flat}\kern-1.00006ptf(\alpha,\beta)$ be the homogeneous polynomial of degree d2d-2 between parentheses at the right-hand-side of (4.2.1). The line $\langle p^{\prime},p^{\prime\prime}\rangle$ is a bitangent to $S$ if and only if $\vphantom{f}{}^{\flat}\kern-1.00006ptf$ has a multiple root, so the polynomial $R$ we are looking for is merely the discriminant of $\vphantom{f}{}^{\flat}\kern-1.00006ptf$ which is none other, up to sign, than the reduced discriminant of $f$, as noticed in (2.10).
Now, the polynomial $f$ is of the form
$f(\alpha,\beta)=\beta^{2}\left(a_{2}\,\alpha^{d-2}+a_{3}\,\alpha^{d-3}\beta+\cdots a_{d}\,\beta^{d-2}\right)$
where the coefficients $a_{i}$ are homogeneous in $p^{\prime}$ and $p^{\prime\prime}$ of degrees did-i and $i$ respectively. Therefore, we deduce from Example (2.10) (see also Proposition (2.14)) that $\operatorname{redDisc}(f)$ has degrees (d2)(d3)(d-2)(d-3) and (d+2)(d3)(d+2)(d-3) in $p^{\prime}$ and $p^{\prime\prime}$ respectively, which concludes the proof. ∎
(4.3) Remark.
The degree of $\operatorname{redDisc}(f)$ in $p^{\prime\prime}$ may be computed alternatively as follows. The plane curve $C_{p^{\prime}}:=\mathbf{T}_{p^{\prime}}S\cap S$ has in general a double point at $p^{\prime}$, and what we want is the number of lines in $\mathbf{T}_{p^{\prime}}S$ passing through $p^{\prime}$ and tangent to $C_{p^{\prime}}$ at some other point. This is the number of ramification points of the projection of $C_{p^{\prime}}$ from $p^{\prime}$; the latter is a (d2):1(d-2):1 map $\bar{C}_{p^{\prime}}\to\mathbf{P}^{1}$, where $\bar{C}_{p^{\prime}}$ denotes the normalization of $C_{p^{\prime}}$ at $p^{\prime}$, so it follows from the Riemann–Hurwitz formula that the number of ramification points equals
$2g(\bar{C}_{p^{\prime}})-2+2(d-2)=\bigl((d-1)(d-2)-2\bigr)-2+2(d-2)=(d+2)(d-3)$
as required.
We were not able to find, on the other hand, a geometric argument to compute the degree in $p^{\prime}$ in a comparable fashion. We wonder wether there is an explanation to why this degree (d2)(d3)(d-2)(d-3) is so nice, in particular in its role in (4.1.1). It is conceivable that it has something to do with the degree of the dual to a smooth plane curve of degree d2d-2.
Proof of Theorem (4.1).
Let $p\in\mathbf{P}^{3}$ be a general point. The locus of those points $p^{\prime}\in S$ such that $(p^{\prime},p)\in\mathcal{T}_{S}$ (equivalently, $p\in\mathbf{T}_{p^{\prime}}S$) is the apparent boundary $\mathrm{D}^{p}S\cap S$. Among these points $p^{\prime}$, the locus of those $p^{\prime}$ for which the line $\langle p,p^{\prime}\rangle$ is a bitangent to $S$ is cut out by $V(R^{p})$, where $R^{p}=R(\_,p)$ with $R$ the polynomial of Proposition (4.2); it is therefore a complete intersection in $\mathbf{P}^{3}$ of type (d,d1,(d2)(d3))(d,d-1,(d-2)(d-3)). One concludes by observing that there are two points $p^{\prime}$ for every bitangent line to $S$ passing through $p$. ∎
A generalization in arbitrary dimension
The following is a fairly direct generalization of Theorem (4.1); as we will see the proof is almost identical to that of Salmon for $n=2$ described in the previous subsection.
(4.1) Theorem.
Consider an integer $n>3$. Let $V$ be a smooth hypersurface of degree $d\geqslant n+1$ in $\mathbf{P}^{n}$, and $p$ be a general point of $\mathbf{P}^{n}$. The number of lines passing through $p$ and with two contact points with $V$ of respective orders $2$ and n1n-1 is
(4.1.1) $\prod\nolimits_{k=0}^{n}(d-k).$
To be precise, the lines under consideration in the above statement are those lines $L$ in $\mathbf{P}^{n}$ such that the intersection scheme of $L$ with $V$ has the form
$L\cap V=2p_{1}+(n-1)p_{2}+p_{3}+\cdots+p_{d-n+1}.$
For $n=3$ this would be bitangent lines as in Theorem (4.1); the reason why we assume $n>3$ above is because if $n=3$ the number (4.1.1) actually counts ordered pairs of points $(p_{1},p_{2})$ such that $\langle p_{1},p_{2}\rangle$ is a bitangent line, so that the number of bitangent lines is only one half of the number (4.1.1), as we have seen indeed in Theorem (4.1).
[…]
Proof of the theorem.
Consider
$\mathcal{T}_{V}=\{(p^{\prime},p^{\prime\prime})\in\mathbf{P}^{n}\times\mathbf{P}^{n}:\quad p^{\prime}\in V,\quad p^{\prime\prime}\in\mathrm{D}_{p^{\prime}}V\cap\cdots\cap\mathrm{D}_{{p^{\prime}}^{n-2}}V,\quad\text{and}\quad p^{\prime\prime}\neq p^{\prime}\}:$
the pair $(p^{\prime},p^{\prime\prime})$ is in $\mathcal{T}_{V}$ if and only if $\langle p^{\prime},p^{\prime\prime}\rangle$ is a line, with contact of order n1n-1 with $V$ at $p^{\prime}$, see [II, Theorem **]. For $(p^{\prime},p^{\prime\prime})\in\mathcal{T}_{S}$, we consider the homogeneous polynomial in $(\alpha,\beta)$
$f(\alpha,\beta)$ $=F(\alpha p^{\prime}+\beta p^{\prime\prime})$
$=\alpha^{d}F(p^{\prime})+\alpha^{d-1}\beta\mathrm{D}^{p^{\prime\prime}}F(p^{\prime})+\frac{1}{2}\alpha^{d-2}\beta^{2}\mathrm{D}^{{p^{\prime\prime}}^{2}}F(p^{\prime})+\cdots+\beta^{d}F(p^{\prime\prime})$
(4.1.2) $=\beta^{n-1}\Bigl(\frac{1}{(n-1)!}\alpha^{d-n+1}\mathrm{D}^{{p^{\prime\prime}}^{n-1}}F(p^{\prime})+\cdots+\beta^{d-n+1}F(p^{\prime\prime})\Bigr),$
and denote by $\vphantom{f}{}^{\flat}\kern-1.00006ptf(\alpha,\beta)$ the degree dn+1d-n+1 polynomial between parentheses in the last expression. The line $\langle p^{\prime},p^{\prime\prime}\rangle$ has an additional tangent point with $V$ if and only if the discriminant of $\vphantom{f}{}^{\flat}\kern-1.00006ptf$ vanishes. For all $i=n-1,\ldots,d$, $\mathrm{D}^{{p^{\prime\prime}}^{i}}F(p^{\prime})$ has degree $i$ in $p^{\prime\prime}$ and did-i in $p^{\prime}$. Therefore, by Example (2.10), $\mathrm{Disc}(\vphantom{f}^{\flat}\kern-1.00006ptf)$ has degree (dn+1)(dn)(d-n+1)(d-n) in $p^{\prime}$.
Now, the locus of those points $p^{\prime}\in V$ such that $(p^{\prime},p)\in\mathcal{T}_{V}$, i.e., such that the line $\langle p,p^{\prime}\rangle$ has contact of order n1n-1 with $V$ at $p^{\prime}$, is the intersection $V\cap\mathrm{D}^{p}V\cap\cdots\cap\mathrm{D}^{p^{n-2}}$ (see [II, Theorem **]), which is the complete intersection of n1n-1 hypersurfaces of degrees $d,d-1,\ldots,d-n+2$ respectively. By the first part of the proof, those $p^{\prime}$ such that the line $\langle p,p^{\prime}\rangle$ has an additional contact of order $2$ are cut out in this complete intersection by an $n$-th equation, of degree (dn+1)(dn)(d-n+1)(d-n). One thus gets a 0-dimensional complete intersection of degree
$d\,(d-1)\cdots(d-n+2)\bigl[(d-n+1)(d-n)\bigr],$
which is isomorphic to the family of lines we are interested in. ∎
A dual generalization in arbitrary dimension
The family of codimension $2$ linear subspaces $\Lambda\subseteq\mathbf{P}^{n}$ passing through n2n-2 points $p_{1},\ldots,p_{n-2}$ in linear general position is isomorphic to $\mathbf{P}^{2}$: indeed, such a linear subspace $\Lambda$ is uniquely determined by the datum of a point in the $2$-dimensional projective quotient
$\mathbf{P}^{n}/\langle p_{1},\ldots,p_{n-2}\rangle$
(i.e., $\mathbf{P}\bigl(\mathbf{k}^{n+1}/\langle\hat{p}_{1},\ldots,\hat{p}_{n-2}\rangle$, see footnote 7). Besides, the Grassmannian of codimension $2$ linear subspaces in $\mathbf{P}^{n}$ has dimension 2(n1)2(n-1), and in this codimension passing through a given point is given by two independent linear conditions.
In $\mathbf{P}^{n}/\langle p_{1},\ldots,p_{n-2}\rangle\cong\mathbf{P}^{2}$, those $\Lambda$ that are tangent to $V$ at one point are parametrized by a curve $C$, and those that are bitangent to $V$ (i.e., tangent at two points) are finitely many. The degree of $C$ is readily computed; the theorem below gives the number of bitangent codimension $2$ linear subspaces.
In Paragraph (4.2) below, we explain how the degree of $C$ may be found using homogeneity properties of the discriminant; we shall see that the number of bitangent codimension $2$ subspaces involves analogously a homogeneity property of the reduced discriminant.
(4.1) Theorem.
Consider an integer $n\geqslant 3$. Let $V$ be a general hypersurface of degree $d\geqslant 4$ in $\mathbf{P}^{n}$, and $p_{1},\ldots,p_{n-2}$ be points in linear general position in $\mathbf{P}^{n}$. Then the number of codimension $2$ linear subspaces $\Lambda\subseteq\mathbf{P}^{n}$ passing through $p_{1},\ldots,p_{n-2}$ and bitangent to $V$ is
$\frac{1}{2}\,d(d-1)^{n-2}(d-2)\,\frac{d(d-1)^{n-2}-(3n-5)d+6(n-2)}{d-2}.$
The peculiar way in which the formula is written is meant to emphasize that $2$ is a root of the polynomial $d(d-1)^{n-2}-(3n-5)d+6(n-2)$.
The transversality condition that must be satisfied by $V$ for the formula to be valid is the following: $V$ must be smooth and have finitely many bitangent codimension $2$ subspaces through general points $p_{1},\ldots,p_{n-2}$, each being tangent at only finitely many points. If it happens that some bitangent subspace has more than two ordinary tangency points, then it has to be counted with the appropriate multiplicity.
We will not insist on verifying that the open subset of the linear space of all degree $d$ hypersurfaces defined by this transversality condition is indeed non-empty, but we will indicate where it is needed in the proof.
The degree of the curve $C\subseteq\mathbf{P}^{2}\cong\mathbf{P}^{n}/\langle p_{1},\ldots,p_{n-2}\rangle$ of simply tangent codimension $2$ subspaces is the number of its intersection points with a given line in $\mathbf{P}^{2}$; it is also the number of (n2)(n-2)-planes $\Lambda\subseteq\mathbf{P}^{n}$ tangent to $V$, passing through $p_{1},\ldots,p_{n-2}$, and contained in a fixed hyperplane $H$ which itself passes through $p_{1},\ldots,p_{n-2}$.
The degree of $C$ may be computed using the homogeneity properties of the (plain) discriminant. Let $F\in\mathbf{k}[X_{0},\ldots,X_{n}]$ be an equation of $V$. We consider the polynomial $f\in\mathbf{k}[A,A_{1},\ldots,A_{n-2}]$, depending on the (implicit) parameters $p,p_{1},\ldots,p_{n-2}\in\mathbf{P}^{n}$, defined by
$f(A,A_{1},\ldots,A_{n-2})=F(A.p+A_{1}.p_{1}+\cdots+A_{n-2}.p_{n-2})$
(with the usual abuse of notation that we do not distinguish between $p,p_{1},\ldots,p_{n-2}\in\mathbf{P}^{n}$ and their representatives $\hat{p},\hat{p}_{1},\ldots,\hat{p}_{n-2}\in\mathbf{k}^{n+1}$). Then,
$f(A,A_{1},\ldots,A_{n-2})=\\ A^{d}F(p)+A^{d-1}\mathrm{D}^{A_{1}.p_{1}+\cdots+A_{n-2}.p_{n-2}}F(p)+\,\cdots\,+\frac{1}{d!}\mathrm{D}^{(A_{1}.p_{1}+\cdots+A_{n-2}.p_{n-2})^{d}}F(p)$
by the Taylor–Newton formula, and from this expression we see that the coefficient of $f$ in the monomial $A^{\alpha}A_{1}^{\alpha_{1}}\cdots A_{n-2}^{\alpha_{n-2}}$ is homogeneous of degree $\alpha$ in the parameter $p$. Therefore, the discriminant $\mathrm{Disc}(f)$ is homogenous of degree $d(d-1)^{n-2}$ in the parameter $p$, by Proposition (2.6.1).
The hypersurface in $\mathbf{P}^{n}$ defined by this homogeneous polynomial is the cone of vertex $\langle p_{1},\ldots,p_{n-2}\rangle$ swept out by those (n2)(n-2)-planes $\langle p,p_{1},\ldots,p_{n-2}\rangle$ that are tangent to $V$. It projects from $\langle p_{1},\ldots,p_{n-2}\rangle$ to the curve $C$ in $\mathbf{P}^{n}/\langle p_{1},\ldots,p_{n-2}\rangle\cong\mathbf{P}^{2}$, hence $C$ has degree $d(d-1)^{n-2}$.
As we will see in the proof of Theorem (4.1), the tangency points with $V$ of the (n2)(n-2)-planes parametrized by $C$ form the complete intersection curve $V\cap\mathrm{D}^{p_{1}}V\cap\cdots\cap\mathrm{D}^{p_{n-2}}V$, which has degree $d(d-1)^{n-2}$ and projects birationally to $C$ from $\langle p_{1},\ldots,p_{n-2}\rangle$; this gives another way of computing the degree of $C$.
(4.3)Proof of Theorem (4.1)
We continue with the setup introduced in (4.2) above, but now we assume that $p$ sits on $V$ and is such that the (n2)(n-2)-plane $\langle p,p_{1},\ldots,p_{n-2}\rangle$ is tangent to $V$ at $p$ itself; equivalently, $p$ sits on the complete intersection curve $V\cap\mathrm{D}^{p_{1}}V\cap\cdots\cap\mathrm{D}^{p_{n-2}}V$.888this equivalence needs the smoothness of $V$; the fact that this is a complete intersection is equivalent to the fact that tangent subspaces through $p_{1},\ldots,p_{n-2}$ form a curve. Under this assumption, the polynomial $f$ takes the form
$f(A,A_{1},\ldots,A_{n-2})=\\ \frac{1}{2}A^{d-2}\mathrm{D}^{(A_{1}.p_{1}+\cdots+A_{n-2}.p_{n-2})^{2}}F(p)+\,\cdots\,+\frac{1}{d!}\mathrm{D}^{(A_{1}.p_{1}+\cdots+A_{n-2}.p_{n-2})^{d}}F(p),$
and its coefficient in the monomial $A^{\alpha}A_{1}^{\alpha_{1}}\cdots A_{n-2}^{\alpha_{n-2}}$ is still homogeneous of degree $\alpha$ in the parameter $p$, if non-zero.
The (n2)(n-2)-plane $\langle p,p_{1},\ldots,p_{n-2}\rangle$ is bitangent to $V$ if and only if the reduced discriminant $\operatorname{redDisc}_{d}^{2}(f)$ vanishes. Thus, the locus of points $p\in V$ such that $\langle p,p_{1},\ldots,p_{n-2}\rangle$ is tangent to $V$ in two points including $p$ is the complete intersection of the curve $V\cap\mathrm{D}^{p_{1}}V\cap\cdots\cap\mathrm{D}^{p_{n-2}}V$ with the hypersurface defined by $\operatorname{redDisc}_{d}^{2}(f)$ as a polynomial in pp.999again, the fact that this is indeed a complete intersection is equivalent to the fact that bitangent subspaces through $p_{1},\ldots,p_{n-2}$ are finitely many, and each of those is tangent to $V$ at only finitely many points; multiple points in this complete intersection will occur if some bitangent subspace has more than two ordinary tangency points with $V$. The latter is homogeneous of degree $d(d-1)^{n-2}-(3n-5)d+6(n-2)$ by (2.14.1). Therefore, the locus of tangency points of bitangent (n2)(n-2)-planes through $p_{1},\ldots,p_{n-2}$ has degree
$d(d-1)^{n-2}\bigl[d(d-1)^{n-2}-(3n-5)d+6(n-2)\bigr].$
The result follows, since each bitangent (n2)(n-2)-plane is tangent in two points. ∎
The flecnodal polynomial
This is carried out by Salmon in [18, §588], with [18, §473] as a fundamental tool. This has already been revisited in modern standards in [3], and actually extended there to hypersurfaces in a projective space of arbitrary dimension, so we are going to be brief.
(5.1)The problem
Let $S$ be a smooth surface in $\mathbf{P}^{3}$ of degree $d>1$. For a general point $p\in S$, there are two lines having intersection multiplicity at $p$ with $S$ strictly greater than $2$: these are the tangent lines to the two smooth branches at $p$ of the curve $\mathbf{T}_{p}S\cap S$, which has an ordinary double point at $p$; thus these two lines intersect $S$ with multiplicity $3$ at $p$.
We shall see that those points $p\in S$ such that there is a line intersecting $S$ with multiplicity strictly greater than $3$ at $p$ is a curve $Fl(S)$, cut out on $S$ by a polynomial of degree 11d2411d-24. We call this curve (respectively polynomial) the flecnodal curve (respectively polynomial) of $S$.
At a general point $p$ of the flecnodal curve, the section of $S$ by its tangent hyperplane $\mathbf{T}_{p}S$ is a curve with a non-degenerate double point at $p$ (i.e., a double point with tangent cone of maximal rank), such that one of its two local branches has an inflection point at $p$. In general, the tangent line to the latter branch meets $S$ with multiplicity $4$ at $p$. Those points $p\in S$ such that the curve $\mathbf{T}_{p}S\cap S$ has a tacnode (i.e., a double point with local equation $y^{2}=x^{4}$) also belong to the flecnodal curve: they are its intersection points with the Hessian of $S$, and they are cuspidal points of the cuspidal double curve of $S^{\vee}$ (the latter curve parametrizes those hyperplanes that cut out a cuspidal curve on $S$).
The following statement is in the spirit of reduced elimination theory, with the polynomial $R$ (in the notation of the proposition) playing the role of the reduced resultant. Here we are not prescribing individual multiplicities at the point $q$ for the polynomials $F,G,H$ as in the reduced resultant situation, but rather a multiplicity for the ideal they generate; then $R$ vanishes when the subscheme defined by this ideal is larger than what is imposed by the hypothesis, as does $\operatorname{redRes}$ in the reduced resultant situation.
(5.2) Proposition ([18, §473]).
Let $F_{q}(p),G_{q}(p),H_{q}(p)$ be three bi-homogeneous polynomials in $p,q\in\mathbf{P}^{3}$, of bi-degrees $(\lambda,\mu)$, $(\lambda^{\prime},\mu^{\prime})$, $(\lambda^{\prime\prime},\mu^{\prime\prime})$ respectively. We assume that for the general point $q\in\mathbf{P}^{3}$,
$\mathrm{mult}_{q}\bigl(\mathrm{V}(F_{q},G_{q},H_{q})\bigr)=\lambda\lambda^{\prime}\lambda^{\prime\prime}.$
The locus of those $q\in\mathbf{P}^{3}$ such that $\mathrm{V}(F_{q},G_{q},H_{q})$ contains a point in addition to $q$ counted with multiplicity $\lambda\lambda^{\prime}\lambda^{\prime\prime}$ is the zero locus of a homogeneous polynomial $R$ of degree
$\lambda^{\prime}\lambda^{\prime\prime}\mu+\lambda\lambda^{\prime\prime}\mu^{\prime}+\lambda\lambda^{\prime}\mu^{\prime\prime}-\lambda\lambda^{\prime}\lambda^{\prime\prime}.$
Of course, the condition that the scheme $\mathrm{V}(F_{q},G_{q},H_{q})$ contains a point in addition to $q$ is equivalent to its having positive dimension. Salmon claims that it is equivalent to the fact that $\mathrm{V}(F_{q},G_{q},H_{q})$ contains a line; we have not been able to prove it, but this is not needed for the application.
Proof.
We want to characterize when the scheme $\mathrm{V}(F_{q},G_{q},H_{q})$ has positive dimension. The idea is that this is equivalent to its having non-empty intersection with any hyperplane. So let $L$ be a non-zero linear form in $p$, and consider the resultant $\mathrm{Res}(L,F_{q},G_{q},H_{q})$. It follows from the Poisson formula (see, e.g., [3, Prop. 2.2] and the references therein) and our assumption on $F,G,H$ that there exists a polynomial $R$ such that
(5.2.1) $\mathrm{Res}(L,F_{q},G_{q},H_{q})=L(q)^{\lambda\lambda^{\prime}\lambda^{\prime\prime}}\cdot R(L,F_{q},G_{q},H_{q}).$
Computing degrees, one sees that $R$ is homogeneous of degree 0 in the coefficients of $L$, i.e., it does not depend on $L$. It follows that $V(F_{q},G_{q},H_{q})$ has positive dimension if and only if $R(F_{q},G_{q},H_{q})=0$. Eventually, one computes the degrees of $R(F_{q},G_{q},H_{q})$ using the identity (5.2.1). ∎
(5.3) Theorem.
Let $S$ be a smooth surface in $\mathbf{P}^{3}$ of degree $d$. There exists a homogeneous polynomial $Fl$ of degree 11d2411d-24, such that the locus of points $p\in S$ such that there is a line intersecting $S$ with multiplicity at least $4$ in $p$ is cut out on $S$ by $\mathrm{V}(Fl)$.
Proof.
Let $p\in S$. It follows from [II, Thm. **] that there is a line intersecting $S$ with multiplicity at least $4$ in $p$ if and only if the three polar hypersurfaces $\mathrm{D}_{p}S,\mathrm{D}_{p^{2}}S,\mathrm{D}_{p^{3}}S$ (respectively, the tangent plane, the polar quadric, and the polar cubic of $S$ at $p$) have a common point besides $p$, and that this in turn is equivalent to their having a whole line in common.
On the other hand, $\mathrm{D}_{p}S,\mathrm{D}_{p^{2}}S,\mathrm{D}_{p^{3}}S$, which have degrees $1,2,3$ respectively, intersect with multiplicity $6$ at $p$ by [II, Cor. **]. We are therefore in a position to apply Proposition (5.2), with $(\lambda,\mu)$, $(\lambda^{\prime},\mu^{\prime})$, $(\lambda^{\prime\prime},\mu^{\prime\prime})$ equal to (1,d1)(1,d-1), (2,d2)(2,d-2), (3,d3)(3,d-3) respectively. The result follows. ∎